### v6

# A virtual Clebsch-Gordan formula for $\operatorname{SL}_{2,A}$

Abstract.
If $k$ is a field of characteristic zero and if $\operatorname{SL}(n,k)$ is the group of $n\times n$ matrices with coefficients in $k$ and determinant one, it follows $\operatorname{SL}(n,k)$ is ”semi simple” in the sense that any finite dimensional representation decompose into a finite direct sum of irreducible representations. Moreover any surjective map $p:V\rightarrow W$ of representations has a section $s:W\rightarrow V$ with $s\circ p=Id$ where $s$ is a map of representations. We prove that this property no longer holds for the group scheme $G:=\operatorname{SL}_{2,\mathbb{Z}}$ by giving explicit examples of a surjective map of comodules that does not have a section. Hence $G$ is not ”completely reducible” in this sense. We also prove this result for $\operatorname{SL}_{2,\mathbb{Z}[\frac{1}{m}]}$ and for $\operatorname{SL}_{2,\mathbb{Z}_{(p)}}$ where $\mathbb{Z}_{(p)}$ is the localization of $\mathbb{Z}$ at the prime ideal $(p)$. If $V:=\mathbb{Z}\{e_{1},e_{2}\}$ is the standard comodule we may form the symmetric product $\operatorname{S}^{n}(V)$, and over a field of characteristic zero it follows this construction gives all finite dimensional irreducible modules. Over the integers it follows $\operatorname{S}^{n}(V)$ is no longer irreducible as a comodule, hence there is no classification of $\operatorname{SL}_{2,\mathbb{Z}}$-modules as in the case of a field of characteristic zero. There is a version of the Clebsch-Gordan formula in the following sense: There is for any pair of integers $n,m$ with $1\leq n\leq m$ a finite filtration $F_{i}\subseteq\operatorname{S}^{n}(V)\otimes\operatorname{S}^{m}(V)$ of comodules, with $F_{i}/F_{i+1}\cong\operatorname{S}^{n+m-2i}(V)$. Hence in the grothendieck group we get an equality of classes - a virtual Clenbsch-Gordan formula:
$[\operatorname{S}^{n}(V)][\operatorname{S}^{m}(V)]=\sum_{i=0}^{n}[\operatorname{S}^{n+m-2i}(V)].$
Hence the Clebsch-Gordan formula holds in the grothendieck group of finite rank comodules on $\operatorname{SL}_{2,\mathbb{Z}}$. The filtration $F_{i}$ pulls back to a filtration $F_{i}\otimes_{\mathbb{Z}}A$ of $\operatorname{S}^{n}(V_{A})\otimes\operatorname{S}^{m}(V_{A})$ for any commutative ring $A$, and it gives a virtual Clebsch-Gordan formula in the grothendieck group of ”finite rank” comodules on $\operatorname{SL}_{2,A}$ for any such $A$. Here $V_{A}:=V\otimes_{\mathbb{Z}}A$.
We introduce the notion ”good filtration” for a torsion free comodule on $\operatorname{SL}_{2,\mathbb{Z}}$ and give non trivial examples of torsion free modules equipped with a good filtration.
Key words and phrases:
group scheme, comodule, symmetric power, good filtration, Clebsch-Gordan filtration, Clebsch-Gordan formula, complete reducibility
1. Introduction
The aim of the paper is to study torsion free finite rank comodules on the group scheme $\operatorname{SL}_{2,A}$ where $A$ is a Dedekind domain. If $A$ is a field of characteristic zero there is a complete classification of finite dimensional irreducbible comodules. For any such comodule $W$ there is a direct sum decomposition $W\cong\oplus_{i=1}^{n}\operatorname{S}^{d_{i}}(V)$ where $V$ is the ”standard comodule” and $\operatorname{S}^{d}(V)$ is the d’th symmetric power of $V$. Moreover for any integer $d\geq 0$ it follows the comodule $\operatorname{S}^{d}(V)$ is an irreducible comodule. These properties no longer hold for a general Dedekind domain such as $\mathbb{Z},\mathbb{Z}_{(p)}$ or $\mathbb{Z}[\frac{1}{m}]$. For a torsion free finite rank comodule $W$ on $\operatorname{SL}_{2,A}$ there is no direct sum decomposition into symmetric powers of the standard comodule $V$ (see Theorem 3.4). Moreover the symmetric powers $\operatorname{S}^{d}(V)$ are no longer irreducible comodules on $\operatorname{SL}_{2,A}$ (see Example 3.7).
We say a torsion free comodule $W$ has a ”good filtration” of comodules iff there is a finite filtration of comodules $0:=F_{n+1}\subsetneq F_{n}\subsetneq\cdots\subsetneq F_{0}:=W$ with the following property: There is an isomorphism of comodules $F_{i}/F_{i+1}\cong\operatorname{S}^{d_{i}}(V)$ with $d_{i}\geq 0$ an integer for each $i$. When $A$ is a field of chcaracteristic zero there is the ”classical Clebsch-Gordan formula”. It is the following formula: Let $1\leq n\leq m$ be integers. There is an isomorphism
$\operatorname{S}^{n}(V)\otimes\operatorname{S}^{m}(V)\cong\oplus_{i=0}^{n}\operatorname{S}^{n+m-2i}(V)$
of comodules on $\operatorname{SL}_{2,A}$. This formula gives many filtrations of the tensor product. One of them is the following filtration: Let $F(k)_{i}:=\oplus_{j=0}^{n}\operatorname{S}^{n+m-2j}(V)$ It follows we get a filtration - the classical Clebsch-Gordan filtration -
$0:=F(k)_{n+1}\subsetneq F(k)_{n}\subseteq F(k)_{1}\subseteq F(k)_{0}:=\operatorname{S}^{n}(V)\otimes\operatorname{S}^{m}(V)$
of modules. The classical Clebsch-Gordan filtration is an example of a good filtration of $\operatorname{S}^{n}(V)\otimes\operatorname{S}^{m}(V)$. In this paper we study the tensor product $\operatorname{S}^{n}(V)\otimes\operatorname{S}^{m}(V)$ when $A$ is a Dedekind domain. There is a ”canonical” filtration called the universal Clebsch-Gordan filtration:
$0:=F_{n+1}\subsetneq F_{n}\subseteq F_{n-1}\subseteq\cdots\subseteq F_{1}\subseteq F_{0}:=\operatorname{S}^{n}(V)\otimes\operatorname{S}^{m}(V).$
This filtration is a filtration of $G$-comodules with $F_{i}/F_{i+1}\cong\operatorname{S}^{n+m-2i}(V)$ for all $i=0,..,n$ (see Corollary 2.9). The filtration $F_{i}$ pulls back to the ”classical Clebsch-Gordan filtration” filtration in the case when $k$ is a field of characteristic zero. It follows the universal Clebsch-Gordan filtration is a good filtration of comodules of $\operatorname{S}^{n}(V)\otimes\operatorname{S}^{m}(V)$ for the group scheme $\operatorname{SL}_{2,A}$ over any commutative ring $A$. Hence the Clebsch-Gordan decomposition does not generalize to an arbitrary commutative ring but the classical Clebsch-Gordan filtration does.
We study the ”complete reducibility property” for group schemes $\operatorname{SL}_{2,A}$ where $A:=\mathbb{Z},\mathbb{Z}_{(p)},\mathbb{Z}[\frac{1}{m}]$, and find that these group schemes are not ”completely reducible” in the sense that there are surjections of comodules $\pi:V\rightarrow W$ that do not have a section (see Theorem 3.4). Moreover: The comodule $\operatorname{S}^{n}(V)$ is no longer irreducible when $A$ is one of the rings $\mathbb{Z},\mathbb{Z}_{(p)}$ or $\mathbb{Z}[\frac{1}{m}]$. Hence there is no classification of irreducible $\operatorname{SL}_{2,A}$-comodules in terms of direct sums of the symmetric powers $\operatorname{S}^{n}(V)$ of the standard module $V$.
We give explicit examples of comodules $W,W^{*}$ on $\operatorname{SL}_{2,\mathbb{Z}}$ with the property that $W$ and $W^{*}$ have the same weight space decomposition but where $W$ and $W^{*}$ are non isomorphic as comodules. There is an isomorphism $W\otimes_{\mathbb{Z}}\mathbb{Z}[\frac{1}{2}]\cong W^{*}\otimes_{\mathbb{Z}}\mathbb{Z}[\frac{1}{2}]$ as comodules on $\operatorname{SL}_{2,\mathbb{Z}[\frac{1}{2}]}$. Hence the modules $W,W^{*}$ are not classified by their weights and weight spaces (see Theorem 5.4). We also give examples of comodules with no good filtration (see Example 5.3).
Note: In a paper from 2020 (see [9]) I initiated a study of flag bundles $\mathbb{F}(E)$ where $E$ is a finite rank vector bundle on $S:=\operatorname{Spec}(A)$ and where $A$ is a Dedekind domain. If $P\subseteq\operatorname{SL}_{n,A}$ is a parabolic subgroup scheme of $\operatorname{SL}_{n,A}$, it follows the quotient $\operatorname{SL}_{n,A}/P$ is such a flag bundle. Any representation $(V,\rho)$ of $P$ with the property that $V$ is a torsion free and finite rank projective $A$-module, gives rise to a finite rank vector bundle $E(\rho)$ on $\operatorname{SL}_{n,A}/P$ and for this reason one wants to be able to classify such representations. It is hoped the constructions introduced in this paper may have applications in the study of such flag bundles and their vector bundles.
2. Symmetric powers and the universal Clebsch-Gordan filtration
In this section we wil study representations of the group scheme $\operatorname{SL}_{2,A}$ where $A$ is any commutative ring and $\operatorname{SL}_{2,\mathbb{Z}}$ where $\mathbb{Z}$ is the ring of integers. We will us the language of ”affine group schemes” introduced in the books of Demazure, Gabriel and Jantzen (see [3] and [8]). If $G:=\operatorname{Spec}(R)$ is an affine group scheme with Hopf algebra $R$ we will let $h_{R}$ denote the ”functor of points” of the group scheme $G$.
Let in the following $A$ be a fixed commutative unital ring and let $V:=A\{e_{1},e_{2}\}$ be the free $A$-module on the elements $e_{1},e_{2}$. Let $x_{11},x_{12},x_{21},x_{22}$ be independent variables over $A$ and let $D:=x_{11}x_{22}-x_{12}x_{21}$ Let $R:=A[x_{ij}]/(D-1)$ and let $\operatorname{SL}_{2,A}:=\operatorname{Spec}(R)$. It follows the functor of points
$h_{R}:\underline{A-alg}\rightarrow\underline{Groups}$
defined by $h_{R}(B):=\operatorname{Hom}_{A-alg}(R,B)$ where $B$ is an $A$-algebra, has the property that $h_{R}(B)$ equals the group of $2\times 2$-matrices with coefficients in $B$ and determinant $1$. Hence the functor $h_{R}$ is an affine group scheme in the sense of Jantzen’s book [8] on algebraic groups. Since $R$ gives rise to an affine group scheme it follows the ring $R$ has the structure as Hopf algebra $(R,\Delta_{R},S,\epsilon)$. Here $\Delta_{R},S,\epsilon$ are maps
$\Delta_{R}:R\rightarrow R\otimes_{A}R,$
$S:R\rightarrow R$
and
$\epsilon:R\rightarrow A$
satisfying a set of criteria. The map $\Delta_{R}$ is the comultiplication, $S$ is the co-inversion and $\epsilon$ is the counit. The map $\Delta_{R}$ is defined as follows
$\Delta_{R}(x_{11}):=x_{11}\otimes x_{11}+x_{12}\otimes x_{21},\Delta_{R}(x_{12}):=x_{11}\otimes x_{12}+x_{12}\otimes x_{22},$
$\Delta_{R}(x_{21}):=x_{21}\otimes x_{11}+x_{22}\otimes x_{21},\Delta_{R}(x_{22}):=x_{21}\otimes x_{12}+x_{22}\otimes x_{22}.$
Moreover
$S(x_{11}):=x_{22},S(x_{12}):=-x_{12},S(x_{21}):=-x_{21},S(x_{22}):=x_{11},$
and
$\epsilon(x_{11}):=\epsilon(x_{22}):=1,\epsilon(x_{12}):=\epsilon(x_{21}):=0.$
One checks the above formulas define a Hopf algebra in the sense of Jantzen’s book. Define the following comodule structure on $V$:
$\Delta_{V}:V\rightarrow R\otimes V$
by
$\Delta_{V}(e_{1}):=x_{11}\otimes e_{1}+x_{12}\otimes e_{2},\Delta_{V}(e_{2}):=x_{21}\otimes e_{1}+x_{22}\otimes e_{2}.$
One checks there are equalitites
$1\otimes\Delta_{V}\circ\Delta_{V}=\Delta_{R}\otimes 1\circ\Delta_{V}$
and
$\epsilon\otimes 1\circ\Delta_{V}=I_{V}.$
Hence $(V,\Delta_{V})$ is a comodule on $\operatorname{SL}_{2,A}$.
Definition 2.1.
Let $(V,\Delta_{V})$ be the standard comodule on $\operatorname{SL}_{2,A}$.
Example 2.2.
A relation to multiplication with a matrix.
The structure $\Delta_{V}$ as comodule on $V$ corresponds to ”matrix multiplication from the right”. Let $v\in V$ be represented by a pair $[a_{1},a_{2}]$ with $a_{i}\in A$, and choose a $A$-valued point $M\in\operatorname{SL}_{2,A}(A)$, where $M$ is a $2\times 2$-matrix with coefficients in $A$ and determinant one. The matrix $M$ acts on the vector $[a_{1},a_{2}]$ using matrix multiplication from the right. There is another obvious comodule structure
$\Delta^{l}_{V}:V\rightarrow V\otimes R$
where we use matrix multiplication from the left.
Example 2.3.
Symmetric powers, exterior powers and ”classical duals”
We may for any integer $n\geq 0$ construct the symmetric power $\operatorname{S}^{n}(V)\cong\oplus_{i=0}^{n}Ae_{1}^{n-i}e_{2}^{i}$. It follows there is a canonical structure as comodule
$\Delta_{n}:\operatorname{S}^{n}(V)\rightarrow R\otimes\operatorname{S}^{n}(V).$
We may take the exterior power $\wedge^{2}V$ and there is a canonical comodule structure
$\Delta_{1}:\wedge^{2}V\rightarrow R\otimes\wedge^{2}V.$
One checks this is the ”trivial” structure. Let $v:=e_{1}\wedge e_{2}$ be a basis for the free rank one $A$-module $\wedge^{2}V$. It follows $\Delta_{1}(av)=1\otimes av$ for any $a\in A$.
Let the ”classical dual” $V^{*}:=A\{x_{1},x_{2}\}$ be the free $A$-module on the basis $x_{i}:=e_{i}^{*}$ and define the following map:
$\Delta_{V^{*}}:V^{*}\rightarrow R\otimes V^{*}$
by
$\Delta_{V^{*}}(x_{1}):=x_{22}\otimes x_{1}-x_{21}\otimes x_{2},\Delta_{V^{*}}(x_{2}):=-x_{12}\otimes x_{1}+x_{11}\otimes x_{2}.$
One checks $(V^{*},\Delta_{V^{*}})$ is a comodule.
Lemma 2.4.
There is an isomorphism of comodules $\phi:V\rightarrow V^{*}$ defined by $\phi(e_{1}):=x_{2},\phi(e_{2}):=-x_{1}$.
Proof.
One checks the map $\phi$ is a map of comodules which is an isomorphism. The Lemma follows. ∎
Hence for any integer $n\geq 0$ it follows there is an isomorphism $\operatorname{S}^{n}(V)\cong\operatorname{S}^{n}(V^{*})$ of comodules. Over a field $k$ this can be deduced from the fact that the canonical map
$f:V\otimes_{k}V\rightarrow\wedge^{2}V$
defined by $f(u\otimes v):=u\wedge v$ is a nondegenerate pairing and the fact that $\wedge^{2}V\cong T$ is the trivial comodule. This result is by the above argument true for $\operatorname{SL}_{2,A}$ over any commutative ring $A$.
Let in the following $A:=\mathbb{Z}$ be the ring of integers. We let $\operatorname{S}^{*}(V):=\oplus_{d\geq 0}\operatorname{S}^{d}(V)\cong\mathbb{Z}[e_{1},e_{2}]$ be the symmetric algebra of $V$ over $\mathbb{Z}$. This is a graded commutative ring with graded pieces $\operatorname{S}^{d}(V)$ homogeneous polynomials in $e_{1},e_{2}$ of degree $d$.
Since $G:=\operatorname{SL}_{2,\mathbb{Z}}$ is a group scheme over $\mathbb{Z}$ it follows its algebra of distributions $\operatorname{Dist}(G)$ acts on $V$ and $\operatorname{S}^{d}(V)$. There is an inclusion $\operatorname{Lie}(G)\subseteq\operatorname{Dist}(G)$ and $\operatorname{Lie}(G)\cong\mathfrak{sl}(2,\mathbb{Z})$ is the Lie algebra of $2\times 2$-matrices with coefficients in $\mathbb{Z}$ and trace equal to zero. It follows $\operatorname{S}^{d}(V)$ is a module on $\operatorname{Lie}(G)$ and it has by [3] a weight space decomposition for the action of $H$. We let $\operatorname{Lie}(G):=\mathbb{Z}\{x,H,y\}$ be the ”standard generators” for $\operatorname{Lie}(G)$. The elements $x,H,y$ are defined as follows:
$x:=\begin{pmatrix}0&1\\ 0&0\end{pmatrix}.$
$H:=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}.$
and
$y:=\begin{pmatrix}0&0\\ 1&0\end{pmatrix}.$
If $v:=e_{1}^{p}e_{2}^{q}\in\operatorname{S}^{d}(V)$ it follows
$H(v)=(p-q)v.$
If we let $V(d):=\operatorname{S}^{d}(V)$ it follows there is by [3] a direct sum decomposition into eigenspaces for the action of $H$:
$V(d)\cong\oplus_{i=0}^{d}V(d)_{-d+2i}$
where $V(d)_{j}$ is the eigenspace of vectors where $H$ acts by multiplying with $j$. As an example: $H(e_{2}^{d})=-de_{2}^{d}$. Hence the ”vector” $v:=e_{2}^{d}$ has weight equal to $-d$.
There is a sequence of maps of $\mathbb{Z}$-modules
$\operatorname{S}^{n-1}(V)\otimes\operatorname{S}^{m-1}(V)\rightarrow^{f_{n,m}}\operatorname{S}^{n}(V)\otimes\operatorname{S}^{m}(V)\rightarrow^{g_{n,m}}\operatorname{S}^{n+m}(V)$
where $g_{n,m}$ is the multiplication map and $f_{n,m}$ is defined as follows:
$f_{n,m}(e_{1}^{p}e_{2}^{q}\otimes e_{1}^{s}e_{2}^{t}):=e_{1}^{p+1}e_{2}^{q}\otimes e_{1}^{s}e_{2}^{t+1}-e_{1}^{p}e_{2}^{q+1}\otimes e_{1}^{s+1}e_{2}^{t}.$
For $W:=\operatorname{S}^{n+m}(V)$ there is a weight space decomposition $W\cong\oplus_{i=0}^{n+m}W_{-(n+m)+2i}$ and the weight of $v:=e_{1}^{p}e_{2}^{q}$ denoted $wt(v)$ equals $p-q$. It follows
$wt(e_{1}^{i}e_{2}^{n+m-i})=2i-(n+m)$. Let $0\leq p\leq n,0\leq s\leq m$ and let
$v_{k}:=e_{1}^{k}e_{2}^{n-k}\otimes e_{1}^{i-k}e_{2}^{m-i+k}.$
It follows
$f_{n,m}(v_{k})=e_{1}^{i}e_{2}^{m+n-i}.$
Let $v:=e_{1}^{p}e_{2}^{n-p}\otimes e_{1}^{s}e_{2}^{m-s}$. It follows
$f_{n,m}(v)=e_{1}^{p+s}e_{1}^{n+m-(p+s)}.$
It follows $wt(f_{n,m}(v))=wt(f_{n,m}(v_{k}))$ iff $p+s=i$. It follows the pair $(p,s)$ looks as follows:
$(p,s)=\{(i,0),(i-1,1),\ldots,(k,i-k),\ldots,(0,i)\}.$
Since $(q,t)=(n-p,m-s)$ it follows the corresponding 4-tuple $(p,q,s,t)$ must look as follows:
$(p,q,s,t)=(k,n-k,i-k,m-i+k)$
for an integer $0\leq k\leq i$. Hence a vector $v:=e_{1}^{p}e_{2}^{n-p}\otimes e_{1}^{s}e_{2}^{m-s}$ has the property that $f_{n,m}(v)=e_{1}^{i}e_{2}^{n+m-i}$ iff there is an integer $k$ with $0\leq k\leq i$ and with
$v_{k}:=e_{1}^{k}e_{2}^{n-k}\otimes e_{1}^{i-k}e_{2}^{m-i+k}.$
Hence an element $v:=\sum_{j=0}^{i}a_{i}v_{i}$ has $f(v)=0$ iff $a_{0}+a_{1}+\cdots+a_{i}=0$. For each $11\leq k\leq i-$ 1 it follows
$v_{k}-v_{k+1}=-ze_{1}^{k}e_{2}^{n-(k+1)}\otimes e_{1}^{i-(k+1)}e_{2}^{m-(i-k)}$
where $z:=e_{1}\otimes e_{2}-e_{2}\otimes e_{1}$.
Lemma 2.5.
There is an exact sequence of free $\mathbb{Z}$-modules
$0\rightarrow\operatorname{S}^{n-1}(V)\otimes\operatorname{S}^{m-1}(V)\rightarrow^{f_{n,m}}\operatorname{S}^{n}(V)\otimes\operatorname{S}^{m}(V)\rightarrow^{g_{n,m}}\operatorname{S}^{n+m}(V)\rightarrow 0$
where $g_{n,m}$ is the canonical ”multiplication map” and $f_{n,m}$ is defined as follows: Let $e_{1}^{p}e_{2}^{q}\otimes e_{1}^{s}e_{2}^{t}$ be an element with $p+q=n-1,s+t=m-1$. Define $f_{n,m}$ as follows:
$f_{n,m}(e_{1}^{p}e_{2}^{q}\otimes e_{1}^{s}e_{2}^{t}):=e_{1}^{p+1}e_{2}^{q}\otimes e_{1}^{s}e_{2}^{t+1}-e_{1}^{p}e_{2}^{q+1}\otimes e_{1}^{s+1}e_{2}^{t}.$
Proof.
It is clear that $Im(f_{n,m})\subseteq ker(g_{n,m})$. We prove the reverse inclusion. By the above calculation it follows a vector $v:=a_{0}v_{0}+\cdots a_{i}v_{i}$ has the property that $f_{n,m}(v)=0$ iff $a_{0}+\cdots+a_{i}=0$. We may write
$v=a_{0}(v_{0}-v_{1})+(a_{0}+a_{1})(v_{1}-v_{2})+\cdots(a_{0}+\cdots a_{i-1})(v_{i-1}-v_{i})+(a_{0}+\cdots+a_{i})v_{i}=$
$a_{0}(v_{0}-v_{1})+\cdots+(a_{0}+\cdots+a_{i-1})(v_{i-1}-v_{i})$
and it follows $v\in Im(f_{n,m})$ since each vector $v_{k}-v_{k+1}$ is in $Im(f_{n,m})$. The Lemma follows.
Let $\operatorname{S}^{*}(V):=\oplus_{n\geq 0}\operatorname{S}^{n}(V)$ be the symmetric algebra of $V$ over $\mathbb{Z}$. Let $W:=\operatorname{S}^{*}(V)\otimes\operatorname{S}^{*}(V):=\oplus_{d\geq 0}W_{d}$ be the graded ring with grading
$W_{d}:=\oplus_{i+j=d}\operatorname{S}^{i}(V)\otimes\operatorname{S}^{j}(V)$
It follows $\operatorname{S}^{*}(V)\cong\mathbb{Z}[e_{1},e_{2}]$ is the polynomial ring on $e_{1},e_{2}$ and $W:=\mathbb{Z}[e_{1},e_{2}]\otimes_{\mathbb{Z}}\mathbb{Z}[e_{1},e_{2}]\cong\mathbb{Z}[e_{1},e_{2},u_{1},u_{2}]$ is the polynomial ring in 4 independent variables.
Lemma 2.6.
There are comodule structures
$\Delta_{1}:\operatorname{S}^{*}(V)\rightarrow R\otimes\operatorname{S}^{*}(V)$
and
$\Delta_{2}:W\rightarrow R\otimes W$
and the canonical multiplication map $g:W\rightarrow\operatorname{S}^{*}(V)$ is a map of comodules.
Proof.
Define the map $\Delta_{1}$ as follows: Let $v:=e_{1}^{p}e_{2}^{q}\in\operatorname{S}^{n}(V)$ be two elements and define
$\Delta_{1}(v):=\Delta_{V}(e_{1})^{p}\Delta_{V}(e_{2})^{q}$
and for any polynomial $f(e_{1},e_{2})\in\operatorname{S}^{*}(V)$ define
$\Delta_{1}(f):=f(\Delta_{V}(e_{1}),\Delta_{V}(e_{2})).$
By definition $\Delta_{1},1\otimes\Delta_{1}$ and $\Delta_{R}\otimes 1$ are maps of $\mathbb{Z}$-algebras. It follows
$1\otimes\Delta_{1}(\Delta_{1}(v))=$
$1\otimes\Delta_{1}(\Delta_{V}(e_{1})^{p}\Delta_{V}(e_{2})^{q}))=$
$((1\otimes\Delta_{1}(x_{11}\otimes e_{1}+x_{12}\otimes e_{2}))^{p}(1\otimes\Delta_{1}(x_{21}\otimes e_{1}+x_{22}\otimes e_{2}))^{q}=$
$(x_{11}\otimes\Delta_{V}(e_{1})+x_{12}\otimes\Delta_{V}(e_{2}))^{p}(x_{21}\otimes\Delta_{V}(e_{1})+x_{22}\otimes\Delta_{V}(e_{2}))^{q}=$
$(\Delta_{R}(x_{11})\otimes e_{1}+\Delta_{R}(x_{12})\otimes e_{2})^{p}(\Delta_{R}(x_{21}\otimes e_{1}+\Delta_{R}(x_{22})\otimes e_{2})^{q}=$
$(\Delta_{R}\otimes 1(\Delta_{V}(e_{1})))^{p}(\Delta_{R}\otimes 1(\Delta_{V}(e_{2})))^{q}=$
$\Delta_{R}\otimes 1(\Delta_{1}(e_{1})^{p}\Delta_{1}(e_{2})^{q})=\Delta_{R}\otimes 1(\Delta_{1}(e_{1}^{p}e_{2}^{q})).$
It follows using a similar calculation that $\Delta_{1}:\operatorname{S}^{*}(V)\rightarrow R\otimes\operatorname{S}^{*}(V)$ is a map of $\mathbb{Z}$-modules satisfying the axioms of being a comodule and the first claim follows. Note: The map $\Delta_{1}$ is a map of $\mathbb{Z}$-algebras.
In a similar way one checks the tensor product $W:=\operatorname{S}^{*}(V)\otimes_{\mathbb{Z}}\operatorname{S}^{*}(V)$ has a comodule structure defined as follows: Let $v\otimes w\in W$ be an element with
$\Delta_{1}(v):=\sum_{i}x_{i}\otimes u_{i},\Delta_{1}(w):=\sum_{j}y_{j}\otimes v_{j}.$
The differential $\Delta_{2}:W\rightarrow R\otimes W$ is defined by $\Delta_{2}(v\otimes w):=\sum_{i,j}x_{i}y_{j}\otimes u_{i}\otimes v_{j}$. Define a map
$g:W\rightarrow\operatorname{S}^{*}(V)$ by $g(v\otimes w):=vw$. It follows
$1\otimes g(\Delta_{2}(v\otimes w))=\sum_{i,j}x_{i}y_{j}\otimes u_{i}v_{j}$
and
$\Delta_{1}(g(v\otimes w))=\Delta_{1}(vw):=\Delta_{1}(v)\Delta_{1}(w)=\sum_{i,j}x_{i}yj\otimes u_{i}v_{j}$
since $\Delta_{1}$ is a map of rings. It follows $g$ is a map of comodules and the Lemma follows. ∎
By definition the map $\Delta_{1}$ is defined as follows: Let $f(e_{1},e_{2})\in\operatorname{S}^{*}(V)$ be a polynomial. We get
$\Delta_{1}(f):=f(\Delta_{V}(e_{1}),\Delta_{V}(e_{2})).$
Let $W:=\operatorname{S}^{*}(V)\otimes\operatorname{S}^{*}(V)$ and let $v\otimes w\in W$ with $\Delta_{1}(v)=\sum_{i}x_{i}\otimes v_{i},\Delta_{1}(w):=\sum_{j}y_{j}\otimes u_{j}$. It follows by definition
$\Delta_{2}(v\otimes w):=\sum_{i,j}x_{i}y_{j}\otimes v_{i}\otimes u_{j}\in R\otimes W.$
Let $z:=e_{1}\otimes e_{2}-e_{2}\otimes e_{1}\in W$. By definition we have
$\Delta_{V}(e_{1}):=x_{11}\otimes e_{1}+X_{12}\otimes e_{2},\Delta_{V}(e_{2}):=x_{21}\otimes e_{1}+x_{22}\otimes e_{2}.$
It follows
$\phi_{z}(v\otimes w):=z(v\otimes w)=e_{1}v\otimes e_{2}w-e_{2}v\otimes e_{1}w.$
The map $\Delta_{2}$ is defined as follows: There is a canonical map
$\eta:R\otimes\operatorname{S}^{*}(V)\otimes R\otimes\operatorname{S}^{*}(V)\rightarrow R\otimes W$
defined by
$\eta(x\otimes u\otimes y\otimes v):=xy\otimes u\otimes v.$
Let $\rho:=\Delta_{1}\otimes\Delta_{1}$. By definition it follows $\Delta_{2}:=\eta\circ\rho$. We want to calculate $\Delta_{2}(\phi_{z}(v\otimes w))$. We get
$\rho(\phi_{z}(v\otimes w)):=\rho(e_{1}v\otimes e_{2}w-e_{2}v\otimes e_{1}w)=$
$\Delta_{1}(e_{1}v)\otimes\Delta_{1}(e_{2}w)-\Delta_{1}(e_{2}v)\otimes\Delta_{1}(e_{1}w)=$
$\Delta_{V}(e_{1})\Delta_{1}(v)\otimes\Delta_{V}(e_{2})\Delta_{1}(w)-\Delta_{V}(e_{2})\Delta_{1}(v)\otimes\Delta_{V}(e_{1})\Delta_{1}(w).$
It follows
$\rho(e_{1}v\otimes e_{2}w)=\Delta_{1}(e_{1}v)\otimes\Delta_{1}(e_{2}w)=\Delta_{V}(e_{1})\Delta_{1}(v)\otimes\Delta_{V}(e_{2})\Delta_{1}(w)=$
$\sum_{i,j}x_{11}x_{i}\otimes e_{1}u_{i}\otimes x_{21}y_{j}\otimes e_{1}v_{j}+x_{11}x_{i}\otimes e_{1}u_{i}\otimes x_{22}y_{j}\otimes e_{2}v_{j}+$
$x_{12}x_{i}\otimes e_{2}u_{i}\otimes x_{21}y_{j}\otimes e_{1}v_{j}+x_{12}x_{i}\otimes e_{2}u_{i}\otimes x_{22}y_{j}\otimes e_{2}v_{j}.$
It follows
$\eta(\rho(e_{1}v\otimes e_{2}w))=$
$\sum_{i,j}x_{11}x_{21}x_{i}y_{j}\otimes e_{1}u_{i}\otimes e_{1}v_{j}+x_{11}x_{22}x_{i}y_{j}\otimes e_{1}u_{i}\otimes e_{2}v_{j}+$
$x_{12}x_{21}x_{i}y_{j}\otimes e_{2}u_{i}\otimes e_{1}v_{j}+x_{12}x_{22}x_{i}y_{j}\otimes e_{2}u_{i}\otimes e_{2}v_{j}.$
Similarly it follows
$\eta(\rho(e_{2}v\otimes e_{1}w))=$
$\sum_{i,j}x_{11}x_{21}x_{i}y_{j}\otimes e_{1}u_{i}\otimes e_{1}v_{j}+x_{12}x_{21}x_{i}y_{j}\otimes e_{1}u_{i}\otimes e_{2}v_{j}+$
$x_{11}x_{21}x_{i}y_{j}\otimes e_{2}u_{i}\otimes e_{1}v_{j}+x_{12}x_{22}x_{i}y_{j}\otimes e_{2}u_{i}\otimes e_{2}v_{j}.$
It follows
$\Delta_{2}(\phi_{z}(v\otimes w))=$
$\sum_{i,j}x_{i}y_{j}\otimes((e_{1}\otimes e_{2}-e_{2}\otimes e_{1})u_{i}\otimes v_{j})=$
$1\otimes z(\sum_{i,j}x_{i}y_{j}\otimes u_{i}\otimes v_{j})=$
$1\otimes\phi_{z}(\sum_{i,j}x_{i}y_{j}\otimes u_{i}\otimes v_{j})=1\otimes\phi_{z}(\Delta_{2}(v\otimes w)).$
Proposition 2.7.
Let $W:=\operatorname{S}^{*}(V)\otimes\operatorname{S}^{*}(V)$ and let $z:=e_{1}\otimes e_{2}-e_{2}\otimes e_{1}\in W$. Define the map $\phi_{z}:W\rightarrow W$ by $\phi_{z}(u):=zu$. It follows $\phi_{z}$ is a map of comodules.
Proof.
Let $v\otimes w\in W$ with $\Delta_{1}(v)=\sum_{i}x_{i}\otimes v_{i},\Delta_{1}(w):=\sum_{j}y_{j}\otimes u_{j}$. It follows by definition
$\Delta_{2}(v\otimes w):=\sum_{i,j}x_{i}y_{j}\otimes v_{i}\otimes u_{j}\in R\otimes W.$
Let $z:=e_{1}\otimes e_{2}-e_{2}\otimes e_{1}\in W$. By the above calculation it follows $\Delta_{2}(\phi_{z}(v\otimes w))=1\otimes z(\Delta_{2}(v\otimes w))=1\otimes\phi_{z}(\Delta_{2}(v\otimes w))$. The Proposition follows.
Theorem 2.8.
Let $1\leq n\leq m$ be a pair of integers. There is an exact sequence of $\operatorname{SL}_{2,\mathbb{Z}}$-comodules
$0\rightarrow\operatorname{S}^{n-1}(V)\otimes\operatorname{S}^{m-1}(V)\rightarrow^{f_{n,m}}\operatorname{S}^{n}(V)\otimes\operatorname{S}^{m}(V)\rightarrow^{g_{n,m}}\operatorname{S}^{n+m}(V)\rightarrow 0$
Proof.
The proof follows from Lemma 2.5 and Proposition 2.7. If $1\leq n\leq m$ are integers it follows the map $\phi_{z}$ induce a map $\phi^{n,m}_{z}$ at the graded pieces:
$\phi^{n,m}_{z}:W_{n+m-2}\rightarrow W_{n}.$
For the module $\operatorname{S}^{n-1}(V)\otimes\operatorname{S}^{m-1}(V)$ we get the map
$f_{n,m}:\operatorname{S}^{n-1}(V)\otimes\operatorname{S}^{m-1}(V)\rightarrow\operatorname{S}^{n}(V)\otimes\operatorname{S}^{m}(V)$
from Lemma 2.5. The Theorem follows. ∎
Corollary 2.9.
Let $1\leq n\leq m$ be a pair of integers. There is a filtration of $\operatorname{SL}_{2,\mathbb{Z}}$-comodules
$0:=F_{n+1}\subsetneq F_{n}\subseteq F_{n-1}\subseteq\cdots\subseteq F_{1}\subseteq F_{0}:=\operatorname{S}^{n}(V)\otimes\operatorname{S}^{m}(V)$
with isomorphisms $F_{i}/F_{i+1}\cong\operatorname{S}^{n+m-2i}(V)$ for $i=0,..,n$.
Proof.
The Corollary follows from Theorem 2.8.
Definition 2.10.
Let $1\leq n\leq m$ be a pair of integers. The filtration of $\operatorname{SL}_{2,\mathbb{Z}}$-comodules
$0:=F_{n+1}\subsetneq F_{n}\subseteq F_{n-1}\subseteq\cdots\subseteq F_{1}\subseteq F_{0}:=\operatorname{S}^{n}(V)\otimes\operatorname{S}^{m}(V)$
from Corollary 2.9 is the universal Clebsch-Gordan filtration.
Corollary 2.11.
(The virtual Clebsch-Gordan formula) Let $1\leq n\leq m$ be a pair of integers. Let $\operatorname{K}(\operatorname{SL}_{2,\mathbb{Z}})$ be the grothendieck group of finite rank comodules on $\operatorname{SL}_{2,\mathbb{Z}}$. There is an equality
$[\operatorname{S}^{n}(V)][\operatorname{S}^{m}(V)]=\sum_{i=0}^{n}[\operatorname{S}^{n+m-2i}(V)].$
in $\operatorname{K}(\operatorname{SL}_{2,\mathbb{Z}})$.
Proof.
The Corollary follows from Theorem 2.8. ∎
Example 2.12.
The classical Clebsch-Gordan formula.
Let $S.=\operatorname{Spec}(\mathbb{Z})$ and let $\pi:\operatorname{SL}_{2,\mathbb{Z}}:=G\rightarrow S$ be the canonical map. Let $k$ be the field of rational numbers and let $T:=\operatorname{Spec}(k)$. It follows there is an isomorphism $G_{k}:=\pi^{-1}(T)\cong\operatorname{SL}(2,k)$ and we may view $G_{k}$ as the generic fiber of the map $\pi$. The group scheme $G_{k}$ is completely reducible and we may pull back $\operatorname{S}^{n}(V)\otimes\operatorname{S}^{m}(V)$ to $G_{k}$ to get the comodule $\operatorname{S}^{n}(V_{k})\otimes_{k}\operatorname{S}^{m}(V_{k})$ where $V_{k}:=V\otimes_{\mathbb{Z}}k$. The Clebsch-Gordan formula holds for $\operatorname{S}^{n}(V_{k})\otimes\operatorname{S}^{m}(V_{k})$: Let $1\leq n\leq m$ be a pair of integers. There is an isomorphism of $G_{k}$-comodules
$\operatorname{S}^{n}(V_{k})\otimes\operatorname{S}^{m}(V_{k})\cong\oplus_{j=0}^{n}\operatorname{S}^{n+m-2j}(V_{k}).$
We get a filtration of $G_{k}$-comodules - the classical Clebsch-Gordan filtration -
$F(k)_{i}:=\sum_{j=i}^{n}\operatorname{S}^{n+m-2j}(V_{k})$
with $F(k)_{i}/F(k)_{i+1}\cong\operatorname{S}^{n+m-2i}(V_{k})$. The filtration of comodules
$0:=F(k)_{n+1}\subsetneq F(k)_{n}\subseteq F(k)_{1}\subseteq F(k)_{0}:=\operatorname{S}^{n}(V_{k})\otimes\operatorname{S}^{m}(V_{k})$
is the pull back of the filtration $F_{i}\subseteq\operatorname{S}^{n}(V)\otimes\operatorname{S}^{m}(V)$ from Lemma 2.9 in the sense there are isomorphisms of comodules
$\pi^{*}(F_{i})\cong F(k)_{i}$
for all $i=0,..,n$. Hence the Clebsch-Gordan formula does not generalize to $\operatorname{SL}_{2,\mathbb{Z}}$ but the filtration does. There are many different filtrations of the comodule $\operatorname{S}^{n}(V_{k})\otimes\operatorname{S}^{m}(V_{k})$. The filtration $F(k)_{i}$ gives an ordering of the direct summands $\operatorname{S}^{n+m-2i}(V_{k})$ of $\operatorname{S}^{n}(V_{k})\otimes\operatorname{S}^{m}(V_{k})$.
When calculating grothendieck groups it is enough to have a filtration of comodules, hence the virtual Clebsch-Gordan formula may have interest for such calculations.
3. On the ”complete reduciblilty” property for $\operatorname{SL}_{2,\mathbb{Z}}$
Let in this section $p\in\mathbb{Z}$ be a prime number and let $\mathbb{Z}_{(p)}$ be the localization of $\mathbb{Z}$ at $p$. Let $m$ be an integer and let $\mathbb{Z}[\frac{1}{m}]$ be the localization of $\mathbb{Z}$ at the set $S:=\{m^{i}\}$. We prove that the group schemes $\operatorname{SL}_{2,\mathbb{Z}},\operatorname{SL}_{2,\mathbb{Z}_{(p)}}$ and $\operatorname{SL}_{2,\mathbb{Z}[\frac{1}{m}]}$ do not satisfy the ”complete reducibility” property. Over a field of $k$ characteristic zero, any finite dimensional comodule on $\operatorname{SL}(2,k)$ satisfies this property. In order to study torsion free comodules on $\operatorname{SL}_{2,\mathbb{Z}}$ we introduce the notion ”good filtration” and discuss the problem of classifying torsion free finite rank comodules with a good filtration. We give explicit non trivial examples of torsion free finite rank module with a good filtration, hence many nontrivial torsion free comodules has such a filtration, and the notion may therefore have general interest in the study of comodules on a reductive group scheme over a Dedekind domain. One may define the notion ”good filtration” for torsion free comodules on $\operatorname{SL}_{2,A}$ for any Dedekind domain $A$.
Let us first introduce some notions. Let $A$ be a commutative unital ring and let $U:=\operatorname{Spec}(A)$. Let $G/U$ be an affine group scheme over $U$.
Definition 3.1.
We say $G$ is completely reducible iff for any surjective map of $G$-comodules $p:V\rightarrow W$ where $V$ and $W$ are finitely generated $A$-modules, there is a map of $G$-comodules $s:W\rightarrow V$ with $s\circ p=Id_{W}$.
Note: A surjective map of $G$-comodules $p:V\rightarrow W$ has a $G$-comodule section $s:W\rightarrow V$ iff there is an isomorphism of $G$-comodules $V\cong K\oplus W$ where $K:=ker(p)$. We also say $G$ satisfies the complete reducibility property if $G$ is completely reducible.
Theorem 3.2.
Let $k$ be a field of characteristic zero and let $G$ be an affine group scheme of finite type over $k$, with Lie algebra $\operatorname{Lie}(G)$. It follows $\operatorname{Lie}(G)$ is semi simple iff the centre of $G$ is finite and any finite dimensional $G$-comodule $V$ is isomorphic to a direct sum $V\cong V_{1}\oplus\cdots\oplus V_{i}$ where $V_{i}$ is an irreducible $G$-comodule.
Proof.
This is proved in [3], Corollary 2.2, page 260. ∎
Corollary 3.3.
Let $k$ be a field of characteristic zero and let $G$ be an affine group scheme of finite type over $k$. If $\operatorname{Lie}(G)$ is semi simple it follows $G$ is completely reducible.
Proof.
Let $p:V\rightarrow W$ be a surjective map of $G$-comodules. It follows from Theorem 3.2 there are direct sum decompositions
$V\cong V_{1}\oplus\cdots\oplus V_{i},W\cong W_{1}\oplus\cdots\oplus W_{j}$
where each $V_{s},W_{t}$ is irreducible. Since for any $s,t$ it follows $\operatorname{Hom}_{G-comod}(V_{s},W_{t})$ is either zero or $k$ we must have $V\cong K\oplus W$with $K:=ker(p)$. The Corollary follows. ∎
Let $V:=A\{e_{1},e_{2}\}$ and consider the sequence of comodules
$0\rightarrow\operatorname{S}^{n-1}(V)\rightarrow^{f_{0,n}}V\otimes\operatorname{S}^{n}(V)\rightarrow^{\pi_{n}}\operatorname{S}^{n+1}(V)\rightarrow 0$
where $g$ is the canonical multiplication map and where $f(v):=zv$ where $z:=e_{1}\otimes e_{2}-e_{2}\otimes e_{1}$. It follows the sequence is an exact sequence of $\operatorname{SL}_{2,A}$-comodules where $A$ is an arbitrary commutative unital ring.
It follows $V\otimes\operatorname{S}^{n}(V)$ has a basis
$V\otimes\operatorname{S}^{n}(V)\cong A\{e_{i}\otimes e_{1}^{n},e_{i}\otimes e_{1}^{n-1}e_{2},\cdots,e_{i}\otimes e_{2}^{n}\}.$
Let for $11\leq i\leq n-$ 1
$v_{i}:=(e_{1}\otimes e_{2}-e_{2}\otimes e_{1})e_{1}^{n-i}e_{2}^{i-1}:=ze_{1}^{n-i}e_{2}^{i-1}.$
It follows we may chose a basis for $V\otimes\operatorname{S}^{n}(V)$ as follows:
$V\otimes\operatorname{S}^{n}(V)\cong$
$A\{e_{1}\otimes e_{1}^{n},e_{1}\otimes e_{1}^{n-1}e_{2},ze_{1}^{n-1},e_{1}\otimes e_{1}^{n-2}e_{2}^{2},ze_{1}^{n-2}e_{2},\ldots,e_{1}\otimes e_{2}^{n-1},ze_{2}^{n-1},e_{2}\otimes e_{2}^{n}\}.$
A section $s$ of the projection map $\pi_{n}:V\otimes\operatorname{S}^{n}(V)\rightarrow\operatorname{S}^{n+1}(V)$ would have to look as follows:
$s(e_{2}^{n+1})=e_{2}\otimes e_{2}^{n},s(e_{1}e_{2}^{n})=e_{1}\otimes e_{2}^{n}+aze_{2}^{n-1}$
with $a\in A$. The element $x\in\operatorname{Lie}(G)$ acts as follows: $x(e_{2}^{n+1})=(n+1)e_{1}e_{2}^{n}$. Moreover $x(e_{2}\otimes e_{2}^{n})=e_{1}\otimes e_{2}^{n}+ne_{2}\otimes e_{1}e_{2}^{n-1}$. It follows
$xs(e_{2}^{n+1})=e_{1}\otimes e_{2}^{n}+n(e_{2}\otimes e_{1}e_{2}^{n-1}).$
It follows
$n(e_{2}\otimes e_{1}e_{2}^{n-1})=ne_{1}\otimes e_{2}^{n}-nze_{2}^{n-1}$
hence for $s(x(e_{s}^{n+1})=xs(e_{2}^{n+1})$ to hold we must have an equality
$nze_{2}^{n-1}=a(n+1)ze_{2}^{n-1}$
and it follows $a:=\frac{n}{n+1}\in A$. From this it follows $\frac{1}{n+1}\in A$. Hence the projection map $\pi$ does not have a section $s$ if $A=\mathbb{Z}$. Hence the group scheme $\operatorname{SL}_{2,\mathbb{Z}}$ does not satisfy the ”complete reducibility” property. If $(p)\subseteq\mathbb{Z}$ is a prime ideal and if $A:=\mathbb{Z}_{(p)}$ is the localization of $\mathbb{Z}$ at $(p)$, it follows the sequence
$0\rightarrow\operatorname{S}^{p-2}(V)\rightarrow^{f}V\otimes\operatorname{S}^{p-1}(V)\rightarrow^{g}\operatorname{S}^{p}(V)\rightarrow 0$
does not split since $p$ is not a unit in $A$. Hence $\operatorname{SL}_{2,\mathbb{Z}_{(p)}}$ does not satisfy the ”complete reducibility” property. If $G:=\operatorname{SL}_{2,\mathbb{Z}[\frac{1}{m}]}$ there is always an integer $n$ where the projection map $\pi_{n}$ does not split, hence $G$ does not satisfy the ”complete reducibility” property.
Theorem 3.4.
Let $m$ be any positive integer and let $p\in\mathbb{Z}$ be a prime number. The group schemes $\operatorname{SL}_{2,\mathbb{Z}},\operatorname{SL}_{2,\mathbb{Z}_{(p)}}$ and $\operatorname{SL}_{2,\mathbb{Z}[\frac{1}{m}]}$ do not satifsy the complete reducibility property.
Proof.
The proof is given above. ∎
Let $\Delta_{V}:V\rightarrow R\otimes_{\mathbb{Z}}V$ be a comodule that is finitely generated as $\mathbb{Z}$-module and let $T(V)\subseteq V$ be the torsion sub module. There is a direct sum decomposition $V\cong E(V)\oplus T(V)$ where $E(V)\cong\mathbb{Z}^{n}$ is a free $\mathbb{Z}$-module of rank $n$. Let $p\in\mathbb{Z}$ be a prime and let $T(V)_{p}$ be the set of elements $v$ in $T(V)$ with $p^{m}v=0$ for some integer $m\geq 1$. It follows there is a direct sum decomposition
$T(V)\cong\oplus_{p\in J}T(V)_{p}$
where $J:=\{p_{1},..,p_{l}\}\subseteq\mathbb{Z}$ is a finite set of distinct prime numbers.
Proposition 3.5.
Let $I:=(m)\subseteq\mathbb{Z}$ be an ideal. Let $W:=\operatorname{S}^{*}(V)/I\operatorname{S}^{*}(V)$ and let $W_{n}:=\operatorname{S}^{n}(V)/I\operatorname{S}^{n}(V)$. There is a canonical comodule structure $\Delta_{I}:W\rightarrow R\otimes W$ with the property that the canonical projection map $\pi:\operatorname{S}^{*}(V)\rightarrow W$ is a map of comodules. For each integer $n\geq 1$ we get a comodule $\Delta_{I,n}:W_{n}\rightarrow R\otimes W_{n}$ and there is an exact sequence of comodules
$0\rightarrow\operatorname{S}^{n}(V)\rightarrow^{\phi}\operatorname{S}^{n}(V)\rightarrow^{p}W_{n}\rightarrow 0$
where $\phi(v):=mv$. It follows $Im(\phi)\subsetneq\operatorname{S}^{n}(V)$ is a strict sub-comodule and the comodule $\operatorname{S}^{n}(V)$ is not irreducible. Assume $V$ is a comodule that is finitely generated as $\mathbb{Z}$-module with a direct sum decomposition $V\cong\mathbb{Z}^{n}\oplus T(V)$ where $T(V)$ is the torsion submodule. There is an exact sequence of comodules
$0\rightarrow T(V)\rightarrow V\rightarrow V/T(V)\rightarrow 0$
For any prime $p\in J$ it follows $T(V)_{p}\subseteq T(V)$ is a sub comodule. It follows the direct sum $T(V)\cong\oplus_{p\in J}T(V)_{p}$ is a direct sum of comodules. Let $E(V):=V/T(V)$. It follows $V\cong E(V)\oplus T(V)$ is a direct sum of comodules.
Proof.
The proof is an exercise. ∎
If $k$ is a field of characteristic zero and if $V$ is any finite dimensional $\operatorname{SL}(2,k)$-module, it follows there is a direct sum decomposition $V\cong\oplus_{i\in I}\operatorname{S}^{d_{i}}(V)$ of representations. The symmetric powers $\operatorname{S}^{d}(V)$ are irreducible when $V:=k\{e_{1},e_{2}\}$ is the standard $\operatorname{SL}(2,k)$-module. Proposition 3.7 proves this property does not hold for the group scheme $\operatorname{SL}_{2,\mathbb{Z}}$. The symmetric power of the standard module $\operatorname{S}^{d}(V)$ is not irreducible and there is no decomposition of the tensor product $\operatorname{S}^{n}(V)\otimes\operatorname{S}^{m}(V)$ into a direct sum of irreducible modules. By the above example it follows $V\otimes\operatorname{S}^{n}(V)$ is not isomorphic to the direct sum $\operatorname{S}^{n-1}(V)\oplus\operatorname{S}^{n+1}(V)$ as comodules. Hence the ”classification” of $\operatorname{SL}_{2,\mathbb{Z}}$-modules of finite rank over $\mathbb{Z}$ is more complicated than the case of $\operatorname{SL}(2,k)$.
Proposition 3.7 proves that for any comodule $V$ that is finitely generated as $\mathbb{Z}$-module, if we want to ”classify” $V$, we may ”classify” the torsion free part $E(V)$ and the $p$-part $T(V)_{p}$ separately, since there is a direct sum of comodules
$V\cong E(V)\oplus(\oplus_{p\in J}T(V)_{p}).$
Definition 3.6.
Let $V$ be the standard comodule on $\operatorname{SL}_{2,\mathbb{Z}}$ from Definition 2.1 and let $W$ be a torsion free $\operatorname{SL}_{2,\mathbb{Z}}$-comodule which is of finite rank as $\mathbb{Z}$-module. We say $W$ has a good filtration if there is a finite filtration of comodules $F_{i}\subseteq W$ with
$F_{i}/F_{i+1}\cong\operatorname{S}^{d_{i}}(V)$
with $d_{i}\geq 0$ an integer for all $i$.
By Corollary 2.9 the following holds: For any pair of integers $1\leq n\leq m$ it follows the tensor product $\operatorname{S}^{n}(V)\otimes\operatorname{S}^{m}(V)$ has a good filtration. We may ask for a classification of torsion free comodules on $\operatorname{SL}_{2,\mathbb{Z}}$ having a good filtration.
Example 3.7.
An elementary example.
Let $G:=\operatorname{SL}_{2,\mathbb{Z}}$ and let $V:=\mathbb{Z}\{e_{1},e_{2}\}$ be the standard $G$-comodule from Definition 2.1. Define for any integer $m\in\mathbb{Z}$ a map $f_{m}:V\rightarrow V$ by $f_{m}(v):=mv$. Let $I:=(m)\subseteq\mathbb{Z}$ be the ideal generated by $m$ and assume $I\neq(0),(1)$. Consider the following exact sequence pf $\mathbb{Z}$-modules:
(3.7.1) $0\rightarrow V\rightarrow^{f_{m}}V\rightarrow Q_{m}\rightarrow 0,$
with $Q_{m}\cong\mathbb{Z}/(m)\{e_{1},e_{2}\}\neq(0)$. It follows 3.7.1 is an exact sequence of $G$-comodules, hence $Im(f_{m})\subsetneq V$ is a strict sub $G$-comodule. It follows $V$ has an infinite set of strict sub comodules hence $V$ is not an irreducible comodule.
Note: Let $k$ be the field of rational numbers and let $G_{k}:=\operatorname{SL}(2,k)$. By [11] it follows the representation rings $\operatorname{R}(G)\cong\operatorname{R}(G_{k})$ are isomorphic. One wants to interpret the results in this section in terms of this isomorphism.
Let $R:=A[x_{ij}]/(D-1)$ with $\operatorname{SL}_{2,A}:=\operatorname{Spec}(R)$ and let $T:R\rightarrow R$ be the following map:
$T(x_{11}):=x_{11},T(x_{12}):=x_{21},T(x_{21}):=x_{12},T(x_{22}):=x_{22}.$
Hence $T^{2}=Id$ is the idenitity and the map $T$ corresponds to taking the tranpose of a matrix. It follows $T$ is an automorphisms of $A$-algebras. Recall the inversion map $S:R\rightarrow R$:
$S(x_{11}):=x_{22},S(x_{12}):=-x_{12},S(x_{21}):=-x_{21},S(x_{22}):=x_{11}.$
It follows $S$ is an automorphism of $A$-algebras. Let $\tilde{T}:=T\circ S$. The following holds:
Lemma 3.8.
There is an equality of maps $\tilde{T}=T\circ S=S\circ T$. There is an equality $\tilde{T}\otimes\tilde{T}\circ\Delta_{R}=\Delta_{R}\circ\tilde{T}$. Moreover $\tilde{T}\circ\tilde{T}=Id$.
Proof.
As an example: There is an equality
$\tilde{T}\otimes\tilde{T}(\Delta_{R}(x_{11}))=x_{22}\otimes x_{22}+x_{21}\otimes x_{12}=\Delta_{R}(x_{22})=\Delta_{R}(\tilde{T}(x_{11}).$
The Lemma follows. ∎
We are going to study the classical dual of a classical representation of an abstract group.
Example 3.9.
The ”classical” dual of a ”classical” representation.
If $G$ is an abstract group acting from the left on a finite dimensional vector space $V$ over a field $k$, with $\sigma:G\times V\rightarrow V$ the action, we may dualize the action $\sigma$ as follows. Define the classical dual action
$\sigma^{*}:G\times V^{*}\rightarrow V^{*}$
by $\sigma^{*}(g,\psi)(v):=\psi(g^{-1}v)$. There is a canonical action of $G$ on the direct sum $V\oplus V^{*}$ defined by $g(v,\psi):=(gv,g\psi)$ and the trace map $tr:V\oplus V^{*}\rightarrow k$ is invariant with respect to this action. Moreover the canonical map $\phi:V\rightarrow(V^{*})^{*}$ is an isomorphism of $G$-modules. We want to do something similar for comodules.
Example 3.10.
The ”tranpose” of a free comodule.
We can try to ”mimic” the construction of the dual action of a representation of an abstract group and try to do something similar for a representation of a group scheme. For a free comodule $(W,\Delta)$ if finite rank over $A$, there will in general be several comodule structures on the dual $W^{*}$. Now we will define the transpose comodule $\Delta^{\tilde{T}}$ of $W$.
Let $W:=A\{e_{1},..,e_{n}\}$ be the free $A$-module of rank $n$ on the elements $e_{i}$, and let $\Delta:W\rightarrow R\otimes W$ be a comodule structure with $\Delta(e_{i}):=\sum_{j}a(i)_{j}\otimes e_{j}$ for $i=1,..,n$.
Let $W^{*}:=A\{x_{1},..,x_{n}\}$ with $x_{i}:=e_{i}^{*}$ and define the map
$\Delta_{1}:W^{*}\rightarrow R\otimes W^{*}$
by
$\Delta_{1}(x_{i}):=\sum_{j}a(i)_{j}\otimes x_{i}.$
Define $\Delta^{\tilde{T}}:W^{*}\rightarrow R\otimes W^{*}$ by $\Delta^{\tilde{T}}:=\tilde{T}\otimes 1\circ\Delta_{1}$.
Proposition 3.11.
The maps $\Delta_{1},\Delta^{\tilde{T}}$ are comodule structures on $W^{*}$. The canonial isomorphism of $A$-modules $\phi:W\rightarrow(W^{*})^{*}$ is an isomorphism of comodules.
Proof.
One checks $\Delta_{1}$ is a comodule structure on $W^{*}$. Assume
$\Delta_{1}(x_{i})=\sum_{l}a(i)_{l}\otimes x_{l}.$
It follows
$1\otimes\Delta_{1}(\Delta_{1}(x_{i}))=\sum_{l,k}a(i)_{l}\otimes a(l)_{k}x_{k}.$
Assume
$\Delta_{R}(a(i)_{l})=\sum_{p}u(i,l)_{p}\otimes v(i,l)_{p}.$
It follows there is an equality
$1\otimes\Delta_{1}(\Delta_{1}(x_{i}))=\sum_{l,k}a(i)_{l}\otimes a(l)_{k}\otimes x_{k}=$
$\sum_{l,p}u(i,l)_{p}\otimes v(i,l)_{p}\otimes x_{l}.$
It follows
$1\otimes\Delta^{\tilde{T}}(\Delta^{\tilde{T}}(x_{i}))=\sum_{l,k}\tilde{T}(a(i)_{l})\otimes\tilde{T}(a(l)_{k})\otimes x_{k}=$
$\sum_{l,p}\tilde{T}(u(i,l)_{p})\otimes\tilde{T}(v(i,l)_{p})\otimes x_{l}=$
$\tilde{T}\otimes\tilde{T}\otimes 1(\Delta_{R}\otimes 1(\Delta_{1}(x_{i}))=\Delta_{R}\otimes 1(\tilde{T}\otimes 1(\Delta_{1}(x_{i}))=\Delta_{R}\otimes 1(\Delta_{2}(x_{i}))$
hence $\Delta_{R}\otimes 1\circ\Delta^{\tilde{T}}=1\otimes\Delta^{\tilde{T}}\circ\Delta^{\tilde{T}}$. The rest of the proof is similar and the Proposition is proved. ∎
Definition 3.12.
Let $(W,\Delta)$ be a comodule with $W$ a free $A$-module of rak $n$. We define $(W^{*},\Delta^{\tilde{T}})$ from Proposition 4.12 to be the transpose comodule of $(W,\Delta)$.
Example 3.13.
The transpose of a comodule on $\operatorname{SL}_{2,\mathbb{Z}}$.
As an example: If $G:=\operatorname{SL}_{2,\mathbb{Z}}$ it follow any comodule $W$ that is finitely generated as $\mathbb{Z}$-module, is the direct sum (a direct sum of comodules) $W\cong E(W)\oplus T(W)$ with $E(W)\cong\mathbb{Z}^{n}$ a free rank $n$ module on $\mathbb{Z}$. The module $E(W)$ is the torsion free part of $W$. We may for such modules always define the transpose comodule $W^{*}\cong E(W)^{*}$ since $T(W)^{*}=(0)$ is zero. Hence for any comodule $W$ on $\operatorname{SL}_{2,\mathbb{Z}}$ that is finitely generated as $\mathbb{Z}$-module we may always define the transpose comodule $W^{*}$ using Proposition 4.12.
Example 3.14.
A class of comodules with a good filtration.
Let $k$ be the field of rational numbers and let $\operatorname{SL}(2,k)$ be the pull back of $\operatorname{SL}_{2,\mathbb{Z}}$ to $k$. There is by [11] an isomorphism
$\eta:\operatorname{K}(\operatorname{SL}_{2,\mathbb{Z}})\cong\operatorname{K}(\operatorname{SL}(2,k))$
of abelian groups and $\operatorname{K}(\operatorname{SL}(2,k))$ is generated by the classes $[\operatorname{S}^{d}(V\otimes k)]$ where $V\otimes k$ is the standard comodule on $\operatorname{SL}(2,k)$. The comodule $V\otimes k$ is the pull back of $V$ to $\operatorname{SL}(2,k)$.
Lemma 3.15.
The group $\operatorname{K}(\operatorname{SL}_{2,\mathbb{Z}})$ is generated by the classes $[\operatorname{S}^{d}(V)]$ where $d\geq 0$ is an integer and $(V,\Delta)$ is the standard comodule. There is an equality of classes
$[\operatorname{S}^{d}(V)^{*}]=[\operatorname{S}^{d}(V)]$
for any integer $d\geq 0$.
Proof.
Let $x\in\operatorname{K}(\operatorname{SL}_{2,\mathbb{Z}})$ be any class. It follows $\eta(x)=\sum_{i}n_{i}[\operatorname{S}^{d_{i}}(V\otimes k)]$ for some integers $d_{i}\geq 0$ and $n_{i}\in\mathbb{Z}$. It follows $y:=\sum_{i}n_{i}[\operatorname{S}^{d_{i}}(V)]\in\operatorname{K}(\operatorname{SL}_{2,\mathbb{Z}})$ and since $\eta$ is injective and $\eta(x-y)=0$, it follows $x=y$. The second statement follows using a simlar argument since there is an isomorphism
$\operatorname{S}^{d}(V)^{*}\otimes_{\mathbb{Z}}k\cong\operatorname{S}^{d}(V\otimes_{\mathbb{Z}}k)$
of representations of $\operatorname{SL}(2,k)$. ∎
Note: Lemma 3.15 and Corollary 2.9 suggests that many torsion free comodules on $\operatorname{SL}_{2,\mathbb{Z}}$ have a good filtration. Since the group scheme $\operatorname{SL}_{2,A}$ fails to be completely reducible for a general Dedekind domain, the existence of a good filtration for a large class of torsion free comodules may have interesting consequences for this study. In the following we give more nontrivial examples of torsion free comodules with a good filtration.
Let $\phi:V\rightarrow V$ be defined by $\phi(e_{1}):=e_{2},\phi(e_{2}):=-e_{1}$. Let $\operatorname{S}^{n}(V):=\mathbb{Z}\{v_{0},..,v_{n}\}$ with $v_{i}:=e_{1}^{n-i}e_{2}^{i}$ and let $x_{i}:=v_{i}^{*}$ be the dual basis. Let $\operatorname{S}^{n}(\phi):=\phi^{n}$ and $\operatorname{S}^{*}(\phi):=\phi^{*}$.
Let
$\tilde{T}\otimes\phi^{*}:R\otimes\operatorname{S}^{*}(V)\rightarrow R\otimes\operatorname{S}^{*}(V)$
be defined by $\tilde{T}\otimes\phi^{*}(x\otimes v):=\tilde{T}(x)\otimes\phi^{*}(v)$. Let $(V,\Delta)$ be the standard comodule, let $(\operatorname{S}^{n}(V),\Delta_{n})$ be the $n$’th symmetric product and let $\operatorname{S}^{n}(V)^{*},\Delta^{\tilde{T}}_{n})$ be the tranpose as defined in Proposition 4.12.
It follows
$\tilde{T}\otimes\phi^{*}(\Delta(e_{1}))=\Delta(e_{2}),\tilde{T}\otimes\phi^{*}(\Delta(e_{2}))=-\Delta(e_{1}).$
It follows
$\tilde{T}\otimes\phi^{*}(\Delta_{n}(v_{i}))=\tilde{T}\otimes\phi^{*}(\Delta_{n}(e_{1}^{n-i}e_{2}^{i}))=(-1)^{i}\Delta_{n}(e_{1}^{i}e_{2}^{n-i})=(-1)^{i}\Delta_{n}(v_{n-i}).$
Define the elements $a(i)_{l}\in R$ as follows:
$\Delta_{n}(v_{i}):=a(i)_{0}\otimes v_{0}+a(i)_{1}\otimes v_{1}+\cdots+a(i)_{n}\otimes v_{n}.$
The elements are well defined since $v_{0},..,v_{n}$ is a basis for $R\otimes\operatorname{S}^{n}(V)$ as $R$-module. The elements $a(i)_{l}$ are ”symmetric” with respect to the involution $\tilde{T}$ in a sense expressed in the following Lemma:
Lemma 3.16.
For all $i,l=0,1,..,n$ there is an equality
$\tilde{T}(a(i)_{l})=(-1)^{i+l}a(n-i)_{n-l}.$
Proof.
It follows
$\tilde{T}\otimes\phi^{*}(\Delta_{n}(v_{i}))=$
$\tilde{T}(a(i)_{0})\otimes v_{n}+\cdots+(-1)^{l}\tilde{T}(a(i)_{l})\otimes v_{n-l}+\cdots+(-1)^{n}\tilde{T}(a(i)_{n})\otimes v_{0}=$
$(-1)^{n}\tilde{T}(a(i)_{n})\otimes v_{0}+\cdots+(-1)^{l}\tilde{T}(a(i)_{l})\otimes v_{n-l}+\cdots+\tilde{T}(a(i)_{0})\otimes v_{n}=$
$(-1)^{i}\Delta_{n}(v_{n-i})=$
$(-1)^{i}a(n-i)_{0}\otimes v_{0}+\cdots+(-1)^{i}a(n-i)_{n-l}\otimes v_{n-l}+\cdots+(-1)^{i}a(n-i)_{n}\otimes v_{n}.$
It follows there is for all $i,l=0,..,n$ an equality
$\tilde{T}(a(i)_{l})=(-1)^{i+l}a(n-i)_{n-l}.$
The Lemma follows. ∎
Theorem 3.17.
Let $(V,\Delta)$ be the standard $\operatorname{SL}_{2,\mathbb{Z}}$-comodule and let $W_{1},W_{2}$ be $\operatorname{SL}_{2,\mathbb{Z}}$-comodules on the form $\operatorname{S}^{n}(V),\operatorname{S}^{m}(V^{*})$ or $(\operatorname{S}^{k}(V)^{*},\Delta^{\tilde{T}}_{k})$ for integers $n,m,k\geq 0$. It follows the tensor power $W_{1}\otimes W_{2}$ has a good filtration.
Proof.
There is an isomorphism of comodules $V\cong V^{*}$ hence $\operatorname{S}^{n}(V)\cong\operatorname{S}^{n}(V^{*})$ are isomorphic as comodules. If $v_{i}:=e_{1}^{n-i}e_{2}^{i},x_{i}:=v_{i}^{*}$ are bases for $\operatorname{S}^{n}(V)$ and $\operatorname{S}^{n}(V)^{*}$, there is an isomorphism of $\mathbb{Z}$-modules
$f:\operatorname{S}^{n}(V)\cong\operatorname{S}^{n}(V)^{*}$
defined by $f(v_{i}):=(-1)^{i}x_{n-i}$. By definition
$\Delta_{n}(v_{i})=\sum_{l}a(i)_{l}\otimes v_{l}$
and
$\Delta^{\tilde{T}}_{n}(x_{i}):=\sum_{l}\tilde{T}(a(i)_{l})\otimes x_{l}.$
We get
$1\otimes f(\Delta_{n}(v_{n-i}))=a(n-i)_{0}\otimes x_{n}-a(n-i)_{1}\otimes x_{n-1}+\cdots+$
$(-1)^{n-l}a(n-i)_{n-l}\otimes x_{l}+\cdots+(-1)^{n}a(n-i)_{n}\otimes x_{0}.$
We also get
$\Delta^{\tilde{T}}_{n}(f(v_{n-i}))=(-1)^{n-i}\Delta^{\tilde{T}}_{n}(x_{i})=$
$(-1)^{n-i}\tilde{T}(a(i)_{0})\otimes x_{0}+(-1)^{n-i}\tilde{T}(a(i)_{1})\otimes x_{1}+\cdots+(-1)^{n-i}\tilde{T}(a(i)_{l})\otimes x_{l}+\cdots+$
$(-1)^{n-i}\tilde{T}(a(i)_{n})\otimes x_{n}.$
By Lemma 3.16 there is an equality $\tilde{T}(a(i)_{l})=(-1)^{i+l}a(n-i)_{n-l}$. Hence $\Delta^{\tilde{T}}_{n}(f(v_{i}))=1\otimes f(\Delta_{n}(v_{i}))$ and we have proved $f$ is an isomorphism of comodules. The Theorem follows from Lemma 2.9.
Note: If $k$ is the field of rational numbers and if $V(k):=V\otimes k$ is the standard comodule on $\operatorname{SL}(2,k)$, there is an isomorphism $V(k)\cong V(k)^{*}$ and $\operatorname{S}^{n}(V(k))\cong\operatorname{S}^{n}(V(k)^{*})$. There is a ”canonical” isomorphism of comodules $\tau:\operatorname{S}^{n}(V(k)^{*})\cong\operatorname{S}^{n}(V(k))^{*}$ defined by
$\tau(\phi_{1}\cdots\phi_{n})(v_{1}\cdots v_{n}):=\sum_{\sigma}\phi_{\sigma^{-1}(1)}(v_{1})\cdots\phi_{\sigma^{-1}(n)}(v_{n}),$
hence we get a ”canonical” isomorphism $\tau:\operatorname{S}^{n}(V(k))\cong\operatorname{S}^{n}(V(k))^{*}$.
The isomorphism in Theorem 3.17 differs from $\tau$. There is a map similar to $\tau$ defined for $\operatorname{SL}_{2,\mathbb{Z}}$ but it is no longer an isomorphism of comodules. The isomorphism $\tau$ is constructed in [7], in section B3, page 476.
4. Generalities on duals and left and right comodules on $\operatorname{SL}_{2,\mathbb{Z}}$
I this section we include some general results on left and right comodules on Hopf algebras. We also give an explicit construction of the dual $(V^{*},\Delta^{*})$ of any comodule $(V,\Delta)$ on a Hopf algebra $R$ in the case when $V$ is a free and finite rank module over the base ring $k$. For the group scheme $\operatorname{SL}_{2,\mathbb{Z}}$ any torsion free finitely generated comodule is free, hence we get a functorial definition of the dual of any torsion free and finitely generated comodule on $\operatorname{SL}_{2,\mathbb{Z}}$. Some of the proofs in the section are left to the reader as an exercise. They are standard proofs involving commutative diagrams and a bit of homological algebra.
We use the construction of the dual to give an explicit and elementary construction of a torsion free comodule $W$ on $\operatorname{SL}_{2,\mathbb{Z}}^{op}$ where $W$ and the dual $W^{*}$ have the same weight space decomposition for the induced action of $\operatorname{Lie}(\operatorname{SL}_{2,\mathbb{Z}}^{op})$, but where $W$ and $W^{*}$ are non isomorphic as comodules on $\operatorname{SL}_{2,\mathbb{Z}}^{op}$. Hence a torsion free comodule is not determined by its weight space decomposition as is the case over a field $k$ of characteristic zero.
Note the following: Let $R$ he a Hopf algebra in the sense of Jantzens book. An affine group scheme is a functor
$h_{R}:\underline{A-alg}\rightarrow\underline{Groups}$
where $\underline{A-alg}$ is the category of commutative $A$-algebras and $\underline{Groups}$ is the category of groups. It is defined by $h_{R}(B):=\operatorname{Hom}_{A-alg}(R,B)$. Let $V\cong A\{e_{1},..,e_{n}\}$ be a free rank $n$ module on $A$ and define the following functor:
$\operatorname{GL}(V):\underline{A-alg}\rightarrow\underline{Groups}$
by $\operatorname{GL}(V)(B):=\operatorname{GL}_{B}(B\otimes_{A}V)$. A ”representation” in the language of Jantzen is a natural transformation of functors $\eta:h_{R}\rightarrow\operatorname{GL}(V)$. Given such a natural transformation $\eta$ we may dualize $\eta$ in the following way: For any $A$-algebra $B$ and any element $g\in h_{R}(B)$ we get an element $\eta_{B}(g)^{-1}\in\operatorname{GL}_{B}(B\otimes_{A}V)$ and we may for any element $f\in(B\otimes_{A}V)^{*}\cong V^{*}\otimes_{A}B$ define $\eta^{*}_{B}(g)(f):=f\circ\eta_{B}(g)^{-1}$. We get the ”dual representation”
$\eta^{*}:h_{R}\rightarrow\operatorname{GL}(V^{*}),$
which is proved to be a natural transformation of functors. When doing this we do not ”worry” about left and right representations and if one wants to define the dual directly on the comodule $(V,\Delta)$ corresponding to $V$ one has to do this. The issue of left and right comodules is not mentioned in the books of Demazure, Gabriel and Jantzens and here I will discuss this construction in detail.
Note: The category $\underline{A-alg}$ of all commutative $A$-algebras is not a small category, hence if you have issues with non standard set theory, you may prefer the approach using comodules. Else you can use the definition given in the books of of Demazure, Gabriel and Jantzen. The approach in this section gives an explicit construction of the dual comodule $(V^{*},\Delta^{*})$ of any comodule $(V,\Delta)$ on $\operatorname{SL}_{2,\mathbb{Z}}$ where $V$ is a finitely generated and torsion free $\mathbb{Z}$-module. If you take the ”derivative” of the dual action $\eta^{*}$, you get the dual action of the Lie algebra $\mathfrak{sl}(n,\mathbb{Z})$ on $V^{*}$, but there is a strict inclusion $\operatorname{U}(\mathfrak{sl}(2,\mathbb{Z}))\subsetneq\operatorname{Dist}(\operatorname{SL}_{2,\mathbb{Z}})$. You may not use the Lie algebra or the enveloping algebra to study the isomorphism class of the dual comodule $(V^{*},\Delta^{*})$. Two comodules $V,W$ on $\operatorname{SL}_{2,\mathbb{Z}}$ may be isomorphic as representations of the Lie algebra, but non-isomorphic as comodules.
Let in this section $\operatorname{SL}_{2,\mathbb{Z}}:=\operatorname{Spec}(R)$ where $(R,\Delta_{R},S,\epsilon)$ is the Hopf algebra introduced in section 1. Let $\sigma:R\otimes_{\mathbb{Z}}R\rightarrow R\otimes_{\mathbb{Z}}R$ be defined by $\sigma(a\otimes b):=b\otimes a$ and let $\tilde{\Delta}_{R}:=\sigma\circ\Delta_{R}$. It follows
$\Delta_{R}(x_{11}):=x_{11}\otimes x_{11}+x_{12}\otimes x_{21},\Delta_{R}(x_{12}):=x_{11}\otimes x_{12}+x_{12}\otimes x_{22},$
$\Delta_{R}(x_{21}):=x_{21}\otimes x_{11}+x_{22}\otimes x_{21},\Delta_{R}(x_{22}):=x_{21}\otimes x_{12}+x_{22}\otimes x_{22}.$
and
$\tilde{\Delta}_{R}(x_{11}):=x_{11}\otimes x_{11}+x_{21}\otimes x_{12},\tilde{\Delta}_{R}(x_{12}):=x_{12}\otimes x_{11}+x_{22}\otimes x_{12},$
$\tilde{\Delta}_{R}(x_{21}):=x_{11}\otimes x_{21}+x_{21}\otimes x_{22},\tilde{\Delta}_{R}(x_{22}):=x_{12}\otimes x_{21}+x_{22}\otimes x_{22}.$
One checks $(R,\tilde{\Delta}_{R},S,\epsilon)$ is a Hopf algebra.
Definition 4.1.
The Hopf algebra $(R,\tilde{\Delta}_{R},S,\epsilon)$ is the opposite Hopf algebra of $(R,\Delta_{R},S,\epsilon)$.
Example 4.2.
A representation and its dual.
Let $V:=\mathbb{Z}\{e_{1},e_{2}\}$ and let $B$ be a commutative $\mathbb{Z}$-algebra. Let $V\otimes_{\mathbb{Z}}B\cong\{e_{1}\otimes 1,e_{2}\otimes 1\}B$ and let $C:=\{e_{1}\otimes 1,e_{2}\otimes 1\}$. Let $V^{*}:=\mathbb{Z}\{x_{1},x_{2}\}$ and let $B\otimes_{\mathbb{Z}}V^{*}:=B\{1\otimes x_{1},1\otimes x_{2}\}$. Let $C^{*}:=\{1\otimes x_{1},1\otimes x_{2}\}$.
If $u:=e_{1}\otimes u_{1},+e_{2}\otimes u_{2}$ with $u_{i}\in B$ write
$[u]_{C}:=\begin{pmatrix}u_{1}\\ u_{2}\end{pmatrix}.$
Hence we let $[u]_{C}$ be a ”column vector”. If $f:=v_{1}x_{1}+v_{2}x_{2}$ we write $[f]_{C^{*}}:=[v_{1},v_{2}]$. Hence we let $[f]_{C^{*}}$ be a row vector. There is a pairing $tr:V^{*}\oplus V\rightarrow\mathbb{Z}$ and we have
$tr(f,u):=f(u):=v_{1}u_{2}+v_{2}u_{2}.$
There are two functors
$h_{R},\operatorname{GL}(V):\underline{\mathbb{Z}-alg}\rightarrow\underline{Groups}$
defined by $h_{R}(B):=\operatorname{Hom}_{\mathbb{Z}-alg}(R,B)$ and $\operatorname{GL}(V)(B):=\operatorname{GL}_{B}(V\otimes_{\mathbb{Z}}B)$. There is a natural tranformation $\rho:h_{R}\rightarrow\operatorname{GL}(V)$ defined as follows:
$\rho_{B}:h_{R}(B)\rightarrow\operatorname{GL}_{B}(V\otimes_{\mathbb{Z}}B)$
By definition $V\otimes B\cong B^{2}$. An element $g\in h_{R}(B)$ is by definition a $2\times 2$ matrix $g:=(a_{ij})$ with $det(g)=1$. Hence we may let $g$ act on $V\otimes B$ using matrix multiplication from the left. One checks $\rho$ defines a natural transformation of functors. The natural transformation $\rho$ is by definition a representation of $\operatorname{SL}_{2,\mathbb{Z}}$ on the module $V$. We may calculate its corresponding comodule and one finds that the comodule $\Delta_{V}:V\rightarrow V\otimes R$ is the following:
$\Delta_{V}(e_{1}):=e_{1}\otimes x_{11}+e_{2}\otimes x_{21},\Delta_{V}(e_{2}):=e_{1}\otimes x_{12}+e_{2}\otimes x_{22}.$
We may dualize $\rho$ to a natural transformation
$\rho^{*}:h_{R}\rightarrow\operatorname{GL}(V^{*})$
and by definition $\rho_{B}^{*}:h_{R}(B)\rightarrow\operatorname{GL}_{B}(B\otimes V^{*})$ is defined as follows: Given any element $f\in B\otimes V^{*}\cong(V\otimes B)^{*}$ we let $\rho^{*}_{B}(g)(f):=f\circ\rho_{B}(g^{-1})$. We get a comodule structure $\Delta_{V^{*}}:V^{*}\rightarrow R^{op}\otimes V^{*}$ defined as follows:
$\Delta_{V^{*}}(x_{1}):=x_{22}\otimes x_{1}-x_{12}\otimes x_{2},\Delta_{V^{*}}(x_{2}):=-x_{21}\otimes x_{1}+x_{11}\otimes x_{2}.$
Let $g^{-1}:=(b_{ij})$. The action $\rho_{B}^{*}(g^{-1})(f)$ is given as follows: Multiply the vector $[f]_{C^{*}}:=[v_{1},v_{2}]$ from the right with the matrix $(b_{ij})$. Hence the dual $V^{*}$ of a left $R$-comodule $V$ is a right $R^{op}$-comodule.
Example 4.3.
A involution of an abstract group
If $G$ is an abstract group with multiplication $m:G\times G\rightarrow G$ defined by $m(g,h):=gh$, define the following operation $m^{*}:G\times G\rightarrow G$ by $m^{*}(g,h):=g\bullet h:=hg$. It follows for any $g,h,k\in G$ that
$(g\bullet h)\bullet k=k(hg)=(kh)g=g\bullet(h\bullet k)$
hence the operation $\bullet$ is associative. It follows $(G,\bullet)$ is a group and the map $\phi:(G,m)\rightarrow(G,\bullet)$ defined by $\phi(g):=g^{-1}$ is an isomorphim of groups with the property that $\phi^{2}=Id_{G}$.
We want to define something similar for group schemes.
Lemma 4.4.
The 4-tuple $(R,\tilde{\Delta}_{R},S,\epsilon)$ is a Hopf algebra. The map $S:R\rightarrow R$ sits in a commutative diagram of map of commutative rings
[…]
There is also an equality of maps $\tilde{\Delta}_{R}\circ S=S\otimes S\circ\Delta_{R}$.
Proof.
The proof is an exercise. ∎
We may view the map $S$ as a group scheme version of the involution $\phi$ from Example 4.6. The map $S$ induce a map of group schemes $\phi_{S}:\operatorname{SL}_{2,\mathbb{Z}}\rightarrow\operatorname{SL}_{2,\mathbb{Z}}^{op}$ where $\operatorname{SL}_{2,\mathbb{Z}}^{op}:=\operatorname{Spec}(R)$ with $\tilde{\Delta}_{R}$ as comultiplication. The map $S$ is an isomorphism of Hopf algebras hence $\phi_{S}$ is an isomorphism of group schemes. We may do a similar construction for any affine group scheme $H$.
Definition 4.5.
Let $H$ be a group scheme. Let $\underline{comod}^{H}$ denote the category of comodules $\Delta:V\rightarrow R\otimes V$ with the obvious maps. Let $\underline{comod}_{H}$ denote the category of maps of comodules $\Delta:V\rightarrow V\otimes R$ with the obvious maps. We let $\underline{comod}^{H}$ be the category of right $H$-comodules and we let $\underline{comod}_{H}$ be the category of left $H$-comodules.
Define for any $\mathbb{Z}$-module $V$ the following maps: $\tau_{1}:R\otimes V\rightarrow V\otimes R$ by $\tau_{1}(x\otimes v):=v\otimes x$. Define $\tau_{2}:V\otimes R\rightarrow R\otimes V$ by $\tau_{2}(v\otimes x):=x\otimes v$. It follows $\tau_{1}$ and $\tau_{2}$ are inverses of each other.
Let $\Delta_{V}:V\rightarrow R\otimes V$ be an $H$-comodule and let $\Delta_{W}:W\rightarrow W\otimes R$ be an $H^{op}$-comodule. Let $\Delta_{1}:=\tau_{1}\circ\Delta_{V}$ and let $\Delta_{2}:=\tau_{2}\circ\Delta_{W}$.
Define the maps $F,G$ as follows:
$F(V,\Delta_{V}):=(V,\Delta_{1}),$
where
$\Delta_{1}:V\rightarrow V\otimes R$
and
$G(W,\Delta_{W}):=(W,\Delta_{2})$
where
$\Delta_{2}:W\rightarrow R\otimes W.$
There is a commutative diagram
[…]
Assume for any $v\in V$ we may write
$\Delta_{V}(v):=\sum_{i}a_{i}\otimes v_{i},$
$\Delta_{V}(v_{i}):=\sum_{j}b(i)_{j}\otimes v(i)_{j},$
and
$\Delta_{R}(a_{i}):=\sum_{k}u(i)_{k}\otimes w(i)_{k}.$
Since the diagram commutes there is an equality
$\sum_{i,j}a_{i}\otimes u(i)_{j}\otimes v(i)_{j}=\sum_{i,k}u(i)_{k}\otimes w(i)_{k}\otimes v_{i}.$
We want to prove there is a commutative diagram
[…]
There is a canonical isomorphism $\eta:R\otimes R\otimes V\cong V\otimes R\otimes R$ defined by $\eta(x\otimes y\otimes v):=v\otimes y\otimes x$.
We get
$\Delta_{1}\otimes 1(\Delta_{1}(v))=\sum_{i,j}v(i)_{j}\otimes b(i)_{j}\otimes a_{i}$
and
$1\otimes\tilde{\Delta}_{R}(\Delta_{1}(v))=\sum_{i,k}u(i)_{k}\otimes w(i)_{k}\otimes v_{i},$
and since $\eta$ is an isomorphism the claim follows.
Lemma 4.6.
The maps $F,G$ defined above give rise to functors
$F:\underline{comod}^{H}\rightarrow\underline{comod}_{H^{op}}\text{ and }G:\underline{comod}_{H^{op}}\rightarrow\underline{comod}^{H}.$
There are isomorphisms of functors $F\circ G=Id,G\circ F\cong Id$ hence $F$ and $G$ are equivalences of categories.
Proof.
Using methods similar to the above calculation one checks for any $H$-comodule $(V,\Delta_{V})$ the pair $(V,\Delta_{1})$ is a comodule on $H^{op}$. There is for any map of comodules $\phi$ an induced map $F(\phi)$ of $H^{op}$-comodules and one checks $F$ and $G$ are well defined functors. Given any $H$-comodule $(V,\Delta_{V})$ it follows $(G\circ F)(V,\Delta_{V}))$ is the comodule $(V,\Delta_{3})$ where $\Delta_{3}:=\tau_{2}\circ\tau_{1}\circ\Delta_{V}=\Delta_{V}$, hence $(G\circ F)(V,\Delta_{V})=(V,\Delta_{V})$. The rest of the Lemma is an exercise.
Lemma 4.7.
Let $k$ be any commutative unital ring and let $V$ be a free finite rank $k$-module. Let $V^{*}:=\operatorname{Hom}_{k}(V,k)$. There is an isomorphism of left $R$-modules
$\rho_{1}:R\otimes_{k}V^{*}\cong\operatorname{Hom}_{R}(V\otimes_{k}R,R)$
defined by $\rho_{1}(a\otimes f)(v\otimes b):=af(vb)$, where $a\otimes f\in R\otimes_{k}V^{*},v\otimes b\in V\otimes_{k}R$. There is an isomorphism of right $R$-modules
$\rho_{2}:V^{*}\otimes_{k}R\cong\operatorname{Hom}_{R}(R\otimes_{k}V,R)$
defined by $\rho_{2}(f\otimes a)(b\otimes v):=f(a(bv))$ where $f\otimes a\in V^{*}\otimes_{k}R,b\otimes v\in R\otimes_{k}V$. Let $\psi:V\otimes_{k}R\cong(V\otimes_{k}R)^{**}$ be the canonical iomorphism. We get a canonical isomorphism
$s_{1}:=\rho_{2}^{-1}\circ\rho_{1}^{*}\circ\psi:V\otimes_{k}R\cong V^{**}\otimes_{k}R,$
and there is an equality $s_{1}\cong t\otimes 1$ with $t:V\cong V^{**}$ the canonical isomorphism. We also get a canonical isomorphim $s_{2}:R\otimes_{k}V\cong(R\otimes_{k}V)^{**}\cong R\otimes_{k}V^{**}$ and there is an equality $s_{2}\cong 1\otimes t$.
Proof.
The proof is an exercise. ∎
Let $\Delta_{V}:V\rightarrow V\otimes_{k}R$ where $(R,\Delta_{R},S,\epsilon)$ is a Hopf algebra over $k$ and $V$ is a free finite rank $k$-module. We get an induced map
$\phi:V\otimes_{k}R\rightarrow V\otimes_{k}R$
and an induced dual map
$\phi^{*}:(V\otimes_{k}R)^{*}\rightarrow(V\otimes_{k}R)^{*}$
and an induced map $\Delta_{V,1}:V^{*}\rightarrow(V\otimes_{k}R)^{*}\cong R\otimes_{k}V^{*}$
defined by $\Delta_{V,1}(f):=\phi^{*}(1\otimes f)$ where $\phi^{*}:V^{*}\otimes_{k}R\rightarrow V^{*}\otimes_{k}R$ is the induced map from the isomorphism $\rho_{2}$ from Lemma 4.7.
Proposition 4.8.
Let $H:=\operatorname{Spec}(R)$ and define $F(V,\Delta_{V}):=(V^{*},\Delta_{V,1})$. It follows $F$ defines a functor
$F:\underline{comod}_{H}\rightarrow\underline{comod}^{H}.$
Proof.
The pair $(V,\Delta_{V})$ is a comodule and this implies the map $\Delta_{V}$ form two commutative diagrams. One proves this implies the map $\Delta_{V,1}$ satisfies the commutative diagrams needed for it to be a comodule in $\underline{comod}^{H}$. Hence it follows $(V^{*},\Delta_{V,1})$ is an object in $\underline{comod}^{H}$. One moreover proves that for any map $f:(V,\Delta_{V})\rightarrow(W,\Delta_{W})$ in $\underline{comod}_{H}$ it follows we get a map
$F(f):(W^{*},\Delta_{W,1})\rightarrow(V^{*},\Delta_{V,1})$
in $\underline{comod}^{H}$, and for two composable maps $f,g$ in $\underline{comod}_{H}$ it follows $F(f\circ g)=F(g)\circ F(g)$. We get a well defined functor and the Proposition follows.
Let $\Delta_{V}:V\rightarrow V\otimes R$ be a comodule on $R$ and consider the map
$\Delta_{V,1}:V^{*}\rightarrow R\otimes V^{*}$
from Proposition 4.8. Let $S$ be the involution of $R$ and consider the canonical map $S\otimes 1:R\otimes V^{*}\rightarrow R^{op}\otimes V^{*}$ where $R^{op}$ is the ”opposite” Hopf algebra of $R$. We define a composed map
$\Delta_{V^{*}}:=S\otimes 1\circ\Delta_{V,1}:V^{*}\rightarrow R^{op}\otimes V^{*}.$
Proposition 4.9.
Let $(V,\Delta_{V})$ be a comodule on $R$ where $V$ is a free and finite rank $k$-module and let $H:=\operatorname{Spec}(R)$. Define $D^{LR}(V,\Delta_{V}):=(V^{*},\Delta_{V^{*}})$. The map $D^{LR}(-)$ gives rise to a functor
$D^{LR}_{H}:\underline{comod}_{H}\rightarrow\underline{comod}^{H^{op}}.$
There is similarly a functor
$D^{RL}_{H}:\underline{comod}^{H}\rightarrow\underline{comod}_{H^{op}}.$
Proof.
The construction of $D^{RL}_{H}$ is similar to the construction of $D^{LR}_{H}$. The proof is an exercise. ∎
Example 4.10.
The dual of the standard comodule $(V,\Delta_{V})$ on $\operatorname{SL}_{2,\mathbb{Z}}$.
Let us define the left standard comodule $(V,\Delta^{l}_{V})$ on $\operatorname{SL}_{2,\mathbb{Z}}$ mentioned in Definition 2.1 as the following map:
$\Delta^{l}_{V}:V\rightarrow V\otimes R.$
Let $V:=\mathbb{Z}\{e_{1},e_{2}\}$ and $V^{*}:=\mathbb{Z}\{x_{1},x_{2}\}$ with $x_{i}:=e_{i}^{*}$ the dual coordinates, and define
$\Delta^{l}_{V}(e_{1}):=e_{1}\otimes x_{11}+e_{2}\otimes x_{21},\Delta^{l}_{V}(e_{2}):=e_{1}\otimes x_{12}+e_{2}\otimes x_{22}.$
One checks we get the following: The map $\Delta_{V,1}$ is the following map
$\Delta_{V,1}(x_{1}):=x_{11}\otimes x_{1}+x_{12}\otimes x_{2},\Delta_{V,1}(x_{2}):=x_{21}\otimes x_{1}+x_{22}\otimes x_{2}.$
The involution $S$ gives us the following dual comodule structure:
$\Delta_{V^{*}}:V^{*}\rightarrow R^{op}\otimes V^{*}$
defined by
$\Delta_{V^{*}}(x_{1}):=x_{22}\otimes x_{1}-x_{12}\otimes x_{2},\Delta_{V^{*}}(x_{2}):=-x_{21}\otimes x_{1}+x_{11}\otimes x_{2}.$
We get the same comodule as in Example 4.2.
We may form the tensor product comodule
$\Delta_{2}:V^{*}\otimes_{k}V\rightarrow V^{*}\otimes_{k}V\otimes_{k}R$
and there are trace maps $tr:V^{*}\otimes V\rightarrow\mathbb{Z}$ and $tr\otimes 1:V^{*}\otimes V\otimes R\rightarrow R$ and one checks there is a commutative diagram. Hence with this definition it follows the ”pairing” $tr:V^{*}\otimes V\rightarrow\mathbb{Z}$ is invariant under the coaction of the Hopf algebra. Hence the functor $D^{LR}(-)$ gives a definition of the dual comodule for comodules that are free and of finite rank over $\mathbb{Z}$.
Note: Any torsion free $\mathbb{Z}$-module is free hence we may define the dual comodule for any torsion free comodule. Note also that the ”classical dual” from Example 2.3 differs from the dual from Definition 4.12 in the sense that the dual is a comodule on the opposite group scheme $\operatorname{SL}_{2,\mathbb{Z}}^{op}$. The ”naive dual” is a comodule on $\operatorname{SL}_{2,\mathbb{Z}}$.
Proposition 4.11.
Let $(R,\Delta_{R},S,\epsilon)$ be a Hopf algebra with $H:=\operatorname{Spec}(R)$. Let $(V,\Delta)\in\underline{comod}^{H}$ be a comodule on $R$ with $V$ a finite rank free $k$-module. Let $(V^{**},\Delta^{**}):=(D^{RL}_{H}\circ D^{LR}_{H^{op}})(V,\Delta)$. It follows $V^{**}$ is the double dual of $V$ and the canonical isomorphism $\phi:V\cong V^{**}$ gives an isomorphism of comodules
$(V,\Delta)\cong(V^{**},\Delta^{**}).$
Similarly if $(W,\Delta_{1})\in\underline{comod}_{H}$ is a comodule and if $(W^{**},\Delta_{1}^{**}):=(D^{LR}_{H}\circ D^{RL}_{H^{op}})(W,\Delta_{1})$ it follows there is a canonical isomorphism
$(W,\Delta_{1})\cong(W^{**},\Delta_{1}^{**})$
of comodules. Here $W^{**}$ is the double dual of $W$.
Proof.
Let $\Delta_{1}:V\rightarrow R\otimes V$ be a comodule with $V:=k\{e_{1},..,e_{n}\}$ and $V^{*}:=k\{x_{1},..,x_{n}\}$ with $x_{i}:=e_{i}^{*}$. Let $x_{i}^{*}$ be the dual basis of the basis $\{x_{i}\}$. Assume
$\Delta_{1}(e_{i}):=a_{i1}\otimes e_{1}+a_{i2}\otimes e_{2}+\cdots+a_{in}\otimes e_{n}.$
By construction it follows the map $\Delta_{1}^{*}:V^{*}\rightarrow V^{*}\otimes R^{op}$ is defined as follows:
$\Delta_{1}^{*}(x_{i}):=x_{1}\otimes S(a_{1i})+x_{2}\otimes S(a_{2i})+\cdots+x_{x}\otimes S(a_{ni}).$
It follows we get a comodule $\Delta_{1}^{**}:V^{*}\rightarrow R\otimes V^{**}$ defined by
$\Delta_{1}^{**}(x_{i}^{*}):=S^{2}(a_{i1})\otimes x_{1}^{*}+S^{2}(a_{i2})\otimes x_{2}^{*}+\cdots+S^{2}(a_{in})\otimes x_{n}^{*}=$
$a_{i1}\otimes x_{1}^{*}+a_{i2}\otimes x_{2}^{*}+\cdots+a_{in}\otimes x_{n}^{*}.$
It follows the canonical map $\phi:V\rightarrow V^{**}$ is an isomorphism of comodules, and the first claim of the Proposition follows. The second claim is similar and the Proposition is proved.
Definition 4.12.
Let $(V,\Delta_{V})\in\underline{comod}_{H}$ be a comodule on $R$ and let
$(V^{*},\Delta_{V^{*}}):=D^{LR}_{H}(V,\Delta_{V})$
be the comodule on $R^{op}$ from Proposition 4.9. We let the pair $(V^{*},\Delta_{V^{*}})$ be the right dual comodule of $(V,\Delta_{V})$. Let $(W,\Delta_{W})\in\underline{comod}^{H}$ and define
$(W^{*},\Delta_{W^{*}}):=D^{RL}_{H}(W,\Delta_{W}).$
We let the pair $(W^{*},\Delta_{W^{*}})$ be the left dual comodule of $(W,\Delta_{W})$.
Note: The dual $(V,\Delta_{V^{*}})$ is a comodule on the opposite Hopf algebra $R^{op}$ (or opposite group scheme $H^{op}$).
When we compose the functors $D^{LR}_{H}$ and $D^{RL}_{H}$ we get canonical isomorphisms of comodules, hence it makes sense to use these functors to define the dual comodule. The definition may be used to define the dual for any comodule $(V,\Delta_{V})$ on $\operatorname{SL}_{2,\mathbb{Z}}$ where $V$ is a finitely generated and torsion free $\mathbb{Z}$-module.
Example 4.13.
The dual of the symmetric square.
Let $\operatorname{SL}_{2,\mathbb{Z}}:=\operatorname{Spec}(R)$ and let $\Delta:V\rightarrow R\otimes V$ be the following comodule: Let $V:=\mathbb{Z}\{e_{1},e_{2}\}$ with $B:=\{e_{1},e_{2}\},C:=\{1\otimes e_{1},1\otimes e_{2}\}$ as bases and with the following comodule map:
$\Delta(e_{1}):=x_{11}\otimes e_{1}+x_{12}\otimes e_{2},\Delta(e_{2}):=x_{21}\otimes e_{1}+x_{22}\otimes e_{2}.$
Here we view $V$ as a free rank 2 module on $\mathbb{Z}$ and an element $u\in V$ expressed in the basis $B$ is a ”vector” on the following form: $u:=u_{1}e_{1}+u_{2}e_{2}$ has $[u]_{B}:=[u_{1},u_{2}]$. Let
$M=\begin{pmatrix}x_{11}&x_{12}\\ x_{21}&x_{22}\end{pmatrix}$
It follows
$[\Delta]^{B}_{C}([u]_{B})=[\Delta(u)]_{C}=[u]_{B}M.$
Hence we multiply the vector $[u]_{B}:=[u_{1},u_{2}]$ from the right with the matrix $M$.
We may construct the symmetric power $\Delta_{2}:\operatorname{S}^{2}(V)\rightarrow\operatorname{S}^{2}(V)\otimes R^{op}$ as follows. Let $\operatorname{S}^{2}(V)\cong\mathbb{Z}\{e_{1}^{2},e_{1}e_{2},e_{2}^{2}\}$ with $B_{2}:=\{e_{1}^{2},e_{1}e_{2},e_{2}^{2}\}$ and let $C_{2}:=\{e_{1}^{2}\otimes 1,e_{1}e_{2}\otimes 1,e_{2}^{2}\otimes 1\}$. Define the following matrix:
$M_{2}=\begin{pmatrix}x_{11}^{2}&x_{11}x_{21}&x_{21}^{2}\\ 2x_{11}x_{12}&x_{11}x_{22}+x_{12}x_{21}&2x_{21}x_{22}\\ x_{12}^{2}&x_{12}x_{22}&x_{22}^{2}\end{pmatrix}.$
Write elements $v\in\operatorname{S}^{2}(V)$ as column vectors. It follows
$[\Delta_{2}]^{B_{2}}_{C_{2}}[v]_{B_{2}}=[\Delta_{2}(u)]_{C_{2}}=M_{2}[v]_{B_{2}}.$
Hence we multiply the vector $[v]_{B_{2}}$ from the left with the matrix $M_{2}$.
Using the methods introduced in this section, we may construct the dual of the symmetric power
$\Delta_{3}:\operatorname{S}^{2}(V)^{*}\rightarrow\operatorname{S}^{2}(V)^{*}\otimes R^{op}$
as follows. Let $\operatorname{S}^{2}(V)^{*}\cong\mathbb{Z}\{(e_{1}^{2})^{*},(e_{1}e_{2})^{*},(e_{2}^{2})^{*}\}$ with $B_{3}:=\{(e_{1}^{2})^{*},(e_{1}e_{2})^{*},(e_{2}^{2})^{*}\}$ and let $C_{3}:=\{(e_{1}^{2})^{*}\otimes 1,(e_{1}e_{2})^{*}\otimes 1,(e_{2}^{2})^{*}\otimes 1\}$. Define the following matrix:
$M_{3}=\begin{pmatrix}x_{22}^{2}&-2x_{12}x_{22}&x_{12}^{2}\\ -x_{21}x_{22}&x_{11}x_{22}+x_{12}x_{21}&-x_{11}x_{12}\\ x_{21}^{2}&-2x_{11}x_{21}&x_{11}^{2}\end{pmatrix}.$
Write elements $w\in\operatorname{S}^{2}(V)^{*}$ as column vectors. It follows
$[\Delta_{3}]^{B_{3}}_{C_{3}}[w]_{B_{3}}=[\Delta_{3}(w)]_{C_{3}}=M_{3}[w]_{B_{3}}.$
Hence we multiply the vector $[w]_{B_{3}}$ from the left with the matrix $M_{3}$.
You may check explicitly that the weight space decomposition of $\operatorname{S}^{2}(V)$ and $\operatorname{S}^{2}(V)^{*}$ agree, hence we cannot distinguish these two representations using weights. You may also check explicitly there is no isomorphism of comodules $\phi:\operatorname{S}^{2}(V)\cong\operatorname{S}^{2}(V)^{*}$. Hence the two comodules $\operatorname{S}^{2}(V)$ and $\operatorname{S}^{2}(V)^{*}$ are not isomorphic as comodules on $\operatorname{SL}_{2,\mathbb{Z}}^{op}$. Similar results hold for higher symmetric powers: The comodules $\operatorname{S}^{d}(V)$ and $\operatorname{S}^{d}(V)^{*}$ have the same weights and ranks of weight spaces, but they are not isomorphic as comodules in general.
Let $B:=A[x_{ij}](D-1)$ with $\operatorname{Spec}(B):=\operatorname{SL}_{2,A}$ and let $V:=A\{e_{1},e_{2}\}$ be the following comodule on $\operatorname{SL}_{2,A}$: Let $\Delta:V\rightarrow R\otimes_{A}V$ be defined by
$\Delta(e_{1}):=x_{11}\otimes e_{1}+x_{12}\otimes e_{2},\Delta(e_{2}):=x_{21}\otimes e_{1}+x_{22}\otimes e_{2}.$
Proposition 4.14.
Let $\operatorname{S}^{2}(V)$ be the symmetric square of $V$ and let $\operatorname{S}^{2}(V)^{*}$ be its dual. There is an isomorphism of comodules $\operatorname{S}^{2}(V)\cong\operatorname{S}^{2}(V)^{*}$ iff $2$ is a unit in $B$. Hence the comodules are isomorphic when $A\cong\mathbb{Z}[\frac{1}{2}]$.
Proof.
The proof follows from the above construction and an explicit calculation. ∎
5. The second symmetric tensor of the standard comodule
In this section we prove that the second symmetric tensor $\operatorname{S}_{2}(V)$ and its dual $\operatorname{S}_{2}(V)^{*}$ have the same weights and weight spaces but are non-isomorphic as comodules on $\operatorname{SL}_{2,\mathbb{Z}}$. We also prove they are isomorphic as comodules on $\operatorname{SL}_{2,\mathbb{Z}[\frac{1}{2}]}$. Hence a torsion free finite rank comodule on $\operatorname{SL}_{2,\mathbb{Z}}$ is not uniquely determined by its weights and weight space decomposition.
[…]
Let $\Delta:V\rightarrow V\otimes_{\mathbb{Z}}R$ be the following comodule: Define
$\Delta(e_{1}):=e_{1}\otimes x_{11}+e_{2}\otimes x_{21},\Delta(e_{2}):=e_{1}\otimes x_{12}+e_{2}\otimes x_{22}.$
Example 5.1.
The $d$’th symmetric tensors $\operatorname{S}_{d}(V)$.
Let $d\geq 2$ be an integer. The symmetric group $S_{d}$ on $d$ letters acts on the $d$’th tensor power $V^{\otimes_{\mathbb{Z}}d}$ and we define the $d$’th symmetric tensors $\operatorname{S}_{d}(V)\subseteq V^{\otimes_{\mathbb{Z}}d}$ as the invariants under this action. There is a comodule structure
$\Delta_{d}:\operatorname{S}_{d}(V)\rightarrow\operatorname{S}_{d}(V)\otimes R$
and one may check that $\operatorname{S}_{d}(V)$ and $\operatorname{S}^{d}(V)$ have the same weights and the same ranks of weight spaces.
Note: When studying group schemes and representations of group schemes one prefer to study the symmetric tensors $\operatorname{S}_{d}(V)$ over the symmetric powers $\operatorname{S}^{d}(V)$, since the symmetric tensors have better functorial properties. There is over a field $k$ of characteristic zero an isomorphism of representations $\operatorname{S}_{d}(V\otimes k)\cong\operatorname{S}^{d}(V\otimes k)$. There is an equality of classes $[\operatorname{S}_{d}(V)]=[\operatorname{S}^{d}(V)]$ in $\operatorname{K}(\operatorname{SL}_{2,\mathbb{Z}})$, hence the grothendieck group $\operatorname{K}(\operatorname{SL}_{2,\mathbb{Z}})$ is generated by the classes $[\operatorname{S}_{d}(V)]$ of the symmetric tensors for $d\geq 0$ an integer.
In this section we will study the symmetric tensor $\operatorname{S}_{2}(V)$ and its dual $\operatorname{S}_{2}(V)^{*}$.
Let
$M=\begin{pmatrix}x_{11}&x_{12}\\ x_{21}&x_{22}\end{pmatrix}.$
Write the vector $u:=u_{1}e_{1}+u_{2}e_{2}$ in the basis $C$ as a column vector
$[u]_{C}=\begin{pmatrix}u_{1}\\ u_{2}\end{pmatrix}.$
It follows the map $\Delta$ is the following map:
$[\Delta(u)]_{C_{1}}=M[u]_{C}.$
Hence if we view the map $\Delta$ as a map from $V$ in the basis $C$ to $V\otimes R$ in the basis $C_{1}$ it follows $\Delta$ is given as left multiplication with the matrix $M$. Hence there is an equality
$M=[\Delta]^{C}_{C_{1}}.$
There is an isomorphism $\operatorname{S}_{2}(V)\cong\mathbb{Z}\{e_{11},e_{12},e_{22}\}$ where $e_{11}:=e_{1}\otimes e_{1},e_{12}:=e_{1}\otimes e_{2}+e_{2}\otimes e_{1}$ and $e_{22}:=e_{2}\otimes e_{2}$. Let $E:=\{e_{11},e_{12},e_{22}\}$. Let $\operatorname{S}_{2}(V)\otimes R\cong\{e_{11}\otimes 1,e_{12}\otimes 1,e_{22}\otimes 1\}R$ and let $E_{1}:=\{e_{11}\otimes 1,e_{12}\otimes 1,e_{22}\otimes 1\}$.
We get a comodule structure $\Delta_{2}:\operatorname{S}_{2}(V)\rightarrow\operatorname{S}_{2}(V)\otimes R$ defined as follows:
$\Delta_{2}(e_{11}):=e_{11}\otimes x_{11}^{2}+e_{12}\otimes x_{11}x_{21}+e_{22}\otimes x_{21}^{2},$
$\Delta_{2}(e_{12}):=e_{11}\otimes 2x_{11}x_{12}+e_{12}\otimes(x_{11}x_{22}+x_{12}x_{21})+e_{22}\otimes 2x_{21}x_{22},$
and
$\Delta_{2}(e_{22}):=e_{11}\otimes x_{12}^{2}+e_{12}\otimes x_{12}x_{22}+e_{22}\otimes x_{22}^{2}.$
Let a vector $u:=u_{1}e_{11}+u_{2}e_{12}+u_{3}e_{22}$ be written as a column vector in the basis $E$:
$[u]_{E}=\begin{pmatrix}u_{2}\\ u_{2}\\ u_{3}\end{pmatrix}.$
We may write the map $\Delta_{2}$ in the bases $E,E_{1}$. Let $M$ be the following matrix:
$M:=\begin{pmatrix}x_{11}^{2}&2x_{11}x_{12}&x_{12}^{2}\\ x_{11}x_{21}&x_{11}x_{22}+x_{12}x_{21}&x_{12}x_{22}\\ x_{21}^{2}&2x_{21}x_{22}&x_{22}^{2}\end{pmatrix}.$
It follows there is an equality
$[\Delta_{2}(u)]_{E_{1}}=M[u]_{E}.$
Let $\operatorname{S}_{2}(V)^{*}\cong\mathbb{Z}\{y_{11},y_{12},y_{22}\}$ with $y_{ij}:=e_{ij}^{*}$ the dual basis, and let $E^{*}:=\{y_{ij}\}$. Represent a vector $v:=v_{1}y_{11}+v_{2}y_{12}+v_{3}y{22}$ in the basis $E^{*}$ as a row vector
$[v]_{E^{*}}:=[v_{1},v_{2},v_{3}].$
Let $R^{op}\otimes\operatorname{S}_{2}(V)^{*}\cong R^{op}\{1\otimes y_{ij}\}$ and let $E_{1}^{*}:=\{1\otimes y_{ij}\}$.
Define the following map:
$\Delta_{2}^{*}:\operatorname{S}_{2}(V)^{*}\rightarrow R^{op}\otimes\operatorname{S}_{2}(V)^{*}$
by
$\Delta_{2}^{*}(v):=S\otimes 1\circ\tau_{2}\circ\Delta_{2}(v)$
where
$\tau_{2}:\operatorname{S}_{2}(V)^{*}\otimes R\rightarrow R\otimes\operatorname{S}_{2}(V)^{*}$
is the map defined by $\tau_{2}(v\otimes z):=z\otimes v$ from Lemma 4.6. Define the map $S\otimes 1:R\otimes\operatorname{S}_{2}(V)^{*}\rightarrow R\otimes\operatorname{S}_{2}(V)^{*}$ using the involution $S$ of $R$. From Lemma 4.6 and Proposition 4.9 it follows $(\operatorname{S}_{2}(V)^{*},\Delta_{2}^{*})$ is a comodule on $R^{op}$. Let $M^{*}$ be the following matrix:
$M^{*}:=\begin{pmatrix}x_{22}^{2}&-2x_{12}x_{22}&x_{12}^{2}\\ -x_{21}x_{22}&x_{11}x_{22}+x_{12}x_{21}&-x_{11}x_{12}\\ x_{21}^{2}&-2x_{11}x_{21}&x_{11}^{2}\end{pmatrix}.$
It follows we may describe the map $\Delta_{2}^{*}$ is the bases $E^{*},E^{*}_{1}$ using the matrix $M^{*}$: There is an equality
$[\Delta_{2}^{*}(v)]_{E^{*}_{1}}=[v]_{E^{*}}M^{*}.$
Hence we multiply the vector $[v]_{E^{*}}$ from the right with the matrix $M^{*}$. Hence the map $[\Delta_{2}^{*}]^{E^{*}}_{E_{1}^{*}}$ is given as follows: We multiply a vector with the matrix $M^{*}$ from the right.
We may using Lemma 4.6 apply the involution $\tau_{2}$ to the comodule $\Delta_{2}:\operatorname{S}_{2}(V)\rightarrow\operatorname{S}_{2}(V)\otimes R$ to get a comodule on the opposite Hopf algebra $R^{op}$. Let $\Delta_{2}^{\tau_{2}}:=\tau_{2}\circ\Delta_{2}$. We get a comodule
$\Delta_{2}^{\tau_{2}}:\operatorname{S}_{2}(V)\rightarrow R^{op}\otimes\operatorname{S}_{2}(V).$
If we write a vector $u:=u_{1}e_{11}+u_{2}e_{12}+u_{3}e_{22}$ in the basis $E$, let $[u]_{E}^{tr}:=[u_{1},u_{2},u_{3}]$ denote the transpose vector, which is a row vector. When we use the involution $\tau_{2}$ we take the transpose of matrices. Let $R^{op}\otimes\operatorname{S}_{2}(V)\cong R^{op}\{1\otimes e_{ij}\}$ and let $E_{2}:=\{1\otimes e_{ij}\}$. In the bases $E,E_{2}$ we may write the map $\Delta_{2}^{\tau_{2}}$ as follows:
$[\Delta_{2}^{\tau_{2}}(u)]_{E_{2}}=[u]_{E}^{tr}M^{tr}.$
Here $M^{tr}$ is the transpose of the matrix $M$ and we multiply the row vector $[u]_{E}^{tr}$ from the right with the matrix $M^{tr}$. The matrix $M^{tr}$ is the following matrix:
$M^{tr}:=\begin{pmatrix}x_{11}^{2}&x_{11}x_{21}&x_{21}^{2}\\ 2x_{11}x_{12}&x_{11}x_{22}+x_{12}x_{21}&2x_{21}x_{22}\\ x_{12}^{2}&x_{12}x_{22}&x_{22}^{2}\end{pmatrix}.$
Let $A\in\operatorname{GL}_{\mathbb{Z}}(\mathbb{Z}^{3})$ be a matrix with
$A:=\begin{pmatrix}a_{11}&a_{12}&a_{13}\\ a_{21}&a_{22}&a_{23}\\ a_{31}&a_{32}&a_{33}\end{pmatrix}.$
One checks there is an isomorphism of comodules $\phi:\operatorname{S}_{2}(V)\cong\operatorname{S}_{2}(V)^{*}$ iff there is a matrix $A\in\operatorname{GL}_{\mathbb{Z}}(\mathbb{Z}^{3})$ with $AM^{tr}=M^{*}A$. This gives a system of equations in the ”variables” $a_{ij}$ and one finds $\phi$ exists iff the matrix $A$ has the following form:
$A:=\begin{pmatrix}0&0&-2a\\ 0&a&0\\ -2a&0&0\end{pmatrix}.$
If $a=1$ it follows $det(A)=4$ hence $A$ is invertible iff $2$ is a unit. Hence such a matrix exsists in $\operatorname{GL}_{\mathbb{Z}[\frac{1}{2}]}(\mathbb{Z}[\frac{1}{2}])^{3})$. There is no such matrix in $\operatorname{GL}_{\mathbb{Z}}(\mathbb{Z}^{3})$.
Theorem 5.2.
The comodules $\operatorname{S}_{2}(V)$ and $\operatorname{S}_{2}(V)^{*}$ have the same weights and ranks of weight spaces but are not isomorphic as comodules on $\operatorname{SL}_{2,\mathbb{Z}}$. The pull backs $\operatorname{S}_{2}(V)\otimes_{\mathbb{Z}}\mathbb{Z}[\frac{1}{2]}]$ and $\operatorname{S}_{2}(V)^{*}\otimes_{\mathbb{Z}}\mathbb{Z}[\frac{1}{2}]$ are isomorphic as comodules on $\operatorname{SL}_{2,\mathbb{Z}[\frac{1}{2}]}$.
Proof.
One checks the comodules have the same weights and ranks of weight spaces and an explicit calcualtion proves they are not isomorphic as comodules on $\operatorname{SL}_{2,\mathbb{Z}}$ using the above constructions. One also proves they are isomorphic as comodules on $\operatorname{SL}_{2,\mathbb{Z}[\frac{1}{2}]}$ using an explicit calculation with the matrices $M^{*}$ and $M^{tr}$. The Theorem follows. ∎
Example 5.3.
Comodules on $\operatorname{SL}_{2,\mathbb{Z}}$ and $\operatorname{SL}_{2,\mathbb{Q}}$.
The comodules $\operatorname{S}^{2}(V)^{*}$ and $\operatorname{S}_{2}(V)^{*}$ do not have a good filtration. Hence it is easy to give explicit examples of comodules on $\operatorname{SL}_{2,\mathbb{Z}}$ not having a good filtration.
Let $S:=\operatorname{Spec}(\mathbb{Z})$ and $T:=\operatorname{Spec}(k)$ with $k$ the field of rational numbers. We may view the group scheme $G_{k}:=\operatorname{SL}(2,k)$ as the special fiber of the canonical map $\pi:\operatorname{SL}_{2,\mathbb{Z}}\rightarrow S$ and we may pull back any comodule on $\operatorname{SL}_{2,\mathbb{Z}}$ to $G_{k}$. Let $\tau:G_{k}\rightarrow\operatorname{SL}_{2,\mathbb{Z}}$ be the canonical map. Given any torsion free finite rank comodule $W$ on $\operatorname{SL}_{2,\mathbb{Z}}$ we get a canonical decomposition
$\tau^{*}(W)\cong\oplus_{i=1}^{l}\operatorname{S}^{d_{i}}(k^{2})$
where $k^{2}$ is the standard comodule on $\operatorname{SL}(2,k)$. We would like for each integer $d\geq 0$ a classification of the comodules $W_{d}$ on $\operatorname{SL}_{2,\mathbb{Z}}$ with $\tau^{*}(W_{d})\cong\operatorname{S}^{d}(k^{2})$. The methods introduced above gives a method to classify all such comodules. As an example:
Theorem 5.4.
The set of finite rank torsion free comodules on $\operatorname{SL}_{2,\mathbb{Z}}$ with $\tau^{*}(W)\cong\operatorname{S}^{2}(k^{2})$ are the following modules:
$\operatorname{S}^{2}(V),\operatorname{S}^{2}(V)^{*},\operatorname{S}_{2}(V),\operatorname{S}_{2}(V)^{*}.$
Proof.
This is an explici calculation using the methods introduced in this section. ∎